### v3

# A Systematic Literature Review of Parameter-Efficient Fine-Tuning for Large Code Models

Abstract.
The rise of Artificial Intelligence (AI)-and particularly Large Language Models (LLMs) for code–has reshaped Software Engineering (SE) by enabling the automation of tasks such as code generation, bug detection, and repair. However, these models require significant computational resources for training and fine-tuning, posing challenges for real-world adoption in resource-constrained environments. To address this, the research community has increasingly turned to Parameter-Efficient Fine-Tuning (PEFT)–a class of techniques that enables the adaptation of large models by updating only a small subset of parameters, rather than the entire model. In this Systematic Literature Review (SLR), we examine the growing application of PEFT techniques–across a wide range of software engineering tasks. We analyze how these methods are used to optimize various deep learning (DL) architectures, focusing on their impact on both performance and efficiency. Our study synthesizes findings from 28 peer-reviewed papers, identifying patterns in configuration strategies and adaptation trade-offs. The outcome of this review is a comprehensive taxonomy that categorizes PEFT usage by task type, distinguishing between generative (e.g., Code Summarization) and non-generative (e.g., Code Clone Detection) scenarios. Our findings aim to inform future research and guide the practical deployment of PEFT in sustainable, AI-powered software development.
Our artifacts are publicly available at https://github.com/alvi75/SLR-PEFT
Parameter Efficient Fine Tuning, PEFT, Large Code Models, Software Engineering, Efficient Training, Systematic Literature Review, SLR
These authors contributed equally to this work.
$c$s: Software and its engineering Software creation and management
1. Introduction
The field of Software Engineering (SE) has evolved rapidly, driven by the increasing integration of Artificial Intelligence (AI) into core development practices. This shift has introduced new paradigms of automation, fundamentally reshaping how developers approach complex tasks across the software lifecycle.
At the center of this transformation are Large Language Models (LLMs)—and more specifically, Large Code Models (LCMs). These, are deep learning (DL) models that pushed the boundaries of software engineering automation–taming the intricacies of tasks that were once thought to require exclusively human expertise. Activities such as Bug Repair (Zhang et al., 2022; Joshi et al., 2023; Ahmed et al., 2022; Fan et al., 2023a; Xia et al., 2023b; Pearce et al., 2023; Jin et al., 2023; Wei et al., 2023b; Kong et al., 2024; Weng and Andrzejak, 2023; Fu et al., 2022), Code Summarization and Documentation (Ahmed et al., 2024; Virk et al., 2024; Yun et al., 2024; Ahmed and Devanbu, 2022; Yun et al., 2024; Mastropaolo et al., 2021a, 2024a), Code Completion (Liu et al., 2020; Austin et al., 2021; Guo et al., 2023; Izadi et al., 2022; Ciniselli et al., 2021a, b), Software Testing (Mastropaolo et al., 2022, 2021b; Tufano et al., 2020, 2022a; Schäfer et al., 2023; Ryan et al., 2024), and Code Review (Tufano et al., 2022b; Lu et al., 2023; Li et al., 2022c, d) have increasingly benefited from the integration of LCMs, which are capable of learning rich code semantics and generating context-aware predictions. These models not only streamline traditionally manual and error-prone tasks but also demonstrate strong potential for improving productivity (Tabarsi et al., 2025; Li et al., 2023b) and software quality (Widyasari et al., 2024; Wadhwa et al., 2024) across a wide range of development scenarios (Mastropaolo et al., 2023, 2024c; Lubos et al., 2024; Yang et al., 2024b; Vitale et al., 2023; Mastropaolo et al., 2024b; Watson et al., 2022; Hou et al., 2024).
However, the remarkable performance exhibited by LCMs comes at a significant cost. These models often contain hundreds of millions to billions of parameters, making them exceptionally resource-intensive to train, fine-tune, and deploy. Their sheer scale demands substantial computational power and memory, which in turn raises critical concerns regarding accessibility, scalability, and environmental sustainability—particularly in terms of energy consumption and carbon footprint (Shi et al., 2024). For many organizations and researchers, fully fine-tuning an LCM for each new task is prohibitively costly, creating a pressing need for more efficient adaptation strategies.
To overcome the limitations of full-model fine-tuning, the research community has increasingly embraced Parameter-Efficient Fine-Tuning (PEFT) as a more sustainable and scalable alternative. PEFT techniques are designed to enable model adaptation by modifying only a small portion of a model’s parameters or introducing lightweight, trainable components. Notable approaches in this category include Low-Rank Adaptation (LoRA) (Hu et al., 2022), Adapter Tuning (Wang et al., 2022a), and Prompt Tuning (Lester et al., 2021). Rather than updating the entire set of parameters in a large code model, these methods introduce targeted, task-specific modifications—such as low-rank transformations or adapter modules—that significantly reduce both computational demands and GPU memory usage. These strategies have demonstrated performance on par with, or in some cases superior to, full fine-tuning (Afrin et al., 2025; Hu et al., 2022; Wang et al., 2022a; Lester et al., 2021; Li and Liang, 2021; Dettmers et al., 2023; Zhang et al., 2023a; Pfeiffer et al., 2020), all while requiring only a fraction of the resources.
In the domain of software engineering, PEFT opens the door to continuous and cost-effective model specialization. For instance, tailoring models to specific repositories, coding styles, or project needs–without the overhead associated with training full-scale models. Despite their rapid proliferation and demonstrated success in natural language processing (Han et al., 2024), a comprehensive understanding of how PEFT methods are applied in software engineering remains lacking. Key questions persist: Which PEFT strategies are favored for adapting LCMs in software engineering? – Are certain model architectures or task types better suited for PEFT? What trade-offs exist? etc. In short, the practical effectiveness and limitations of PEFT techniques in the context of code intelligence have yet to be thoroughly examined.
In this paper, we present a Systematic Literature Review (SLR) characterizing generative and non-generative software engineering-related tasks. The research questions we aim to answer, will highlight three different dimensions concerning PEFT-optimization methods for SE-related tasks, these include: RQ1 the types of software engineering tasks that have been targeted using PEFT techniques; RQ2 the types of LLM architecture (e.g., encoder-decoder) and PEFT methods commonly employed in software engineering applications, and finally; RQ3 the impact on task performance and training efficiency of PEFT-based techniques in generative and non-generative SE tasks, when compared against valid full fine-tuning baselines.
By systematically answering these questions, our study aims to bridge the gap in understanding PEFT’s role in optimizing LCMs for software engineering. We provide a comprehensive taxonomy of PEFT applications in SE and distill empirical insights that can guide both future research and practical adoption of these techniques in AI-driven software development environments. This is accomplished through a comprehensive analysis of the current landscape of efficient model adaptation strategies for both generative and non-generative SE tasks, highlighting how PEFT techniques can make the training and deployment of Large Code Models more practical and sustainable in real-world development settings.
The remainder of this paper is organized as follows: Section 2.1 introduces the use of LLMs in software engineering [^1] Note that in this investigation, the terms LLM4SE (Large Language Models for Software Engineering) and LCMs (Large Code Models) are used interchangeably, as both refer to deep learning models tailored for automating software engineering tasks through code understanding and generation., providing an overview of current practices in SE automation and detailing recent advancements at the overlap of Artificial Intelligence (AI) and SE, a field commonly known with the word AI4SE. This section also examines the evolution and variants of LCMs, with a particular focus on their architectural characteristics, which—as we will show—play a key role in determining the choice of suitable PEFT optimization strategies. These foundational elements are further discussed in Section 2.3, which delves into the principles behind PEFT techniques.
We then present the methodology of our study in Section 3, outlining our approach to data collection, paper selection, and classification. Section 4 provides detailed answers to each of our three research questions, offering insights drawn from the systematic review. In Section 6, we discuss the implications of our findings and suggest promising directions for future research on the application of PEFT in SE.
Finally, we address potential threats to the validity of our study in Section 7, and conclude with a summary of our key takeaways and reflections in Section 9. Central to the contributions of this SLR is the taxonomy introduced in Fig. 3, which synthesizes our findings into a structured classification of PEFT applications across generative and non-generative software engineering tasks. This taxonomy not only distills patterns observed throughout the review but also serves as a practical framework for guiding future applications of PEFT. By clarifying the alignment between PEFT strategies and software engineering task types, the taxonomy equips researchers and practitioners with a clearer understanding of the current landscape—ultimately enabling a more confident and precise application as well as design of future parameter-efficient adaptation strategies.
2. Background and Related Work
In this section, we first provide an overview of recent advancements in LCMs and their impact on various software engineering tasks. We then highlight the evolution of PEFT techniques, detailing their methodologies, advantages, and applications in optimizing LLMs.
2.1. Large Code Models
Recent advancements in Large Code Models (LCMs) have significantly accelerated the automation of software engineering tasks, contributing to notable improvements in both efficiency and accuracy across development pipelines (Hou et al., 2024; Niu et al., 2022, 2023). The remarkable capabilities of LCMs are rooted in their large-scale pre-training on extensive and diverse textual and code corpora, as well as the substantial parameter counts that enhance their learning capacity.
While general-purpose LLMs, such as GPT-4 (Achiam et al., 2023) and Claude (Anthropic, nd), exhibit broad applicability across multiple domains, their generality often limits their effectiveness in code-intensive tasks that require domain-specific knowledge. To address this challenge, researchers have introduced specialized LCMs that are explicitly designed to meet the unique demands of software engineering applications. As a result, LCMs have been successfully deployed across a diverse range of SE tasks, including but not limited to Code Generation (Fan et al., 2023b; Li et al., 2022a; Nijkamp et al., 2023), Code Summarization (Gu et al., 2022; Wang et al., 2023a), Code Clone Detection (Svajlenko and Roy, 2016; White et al., 2016), Bug Fixing (Tufano et al., 2020; Chen et al., 2023a), Requirements Analysis (Marques et al., 2024) and Automated Program Repair (Jiang et al., 2021; Mashhadi and Hemmati, 2021). These models build upon the Transformer architecture introduced by Vaswani et al. (Vaswani et al., 2017) and are typically categorized under three primary architectural frameworks: encoder-only, encoder-decoder, and decoder-only models.
Encoder-only models, such as BERT (Devlin, 2018), are based on the Transformer encoder and are frequently adapted for software engineering use cases. Variants such as CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), RoBERTa (Liu et al., 2019), and ALBERT (Lan et al., 2019) have been widely employed in SE tasks. Additionally, specialized models such as BERTOverflow (Tabassum et al., 2020) and CodeRetriever (Li et al., 2022b) have been designed to capture the structural and semantic nuances of source code by introducing tailored pre-training objectives or by integrating program structure representations. For example, CodeBERT employs token prediction to enhance programming language understanding for downstream tasks such as Code Completion and Documentation Generation (Feng et al., 2020). Furthermore, Salza et al. (Salza et al., 2022) evaluated the capacity of encoder-based models (i.e., BERT, RoBERTa) to capture both natural language and source code semantics, ultimately improving Code Search and Retrieval tasks.
Encoder-decoder models, including PLBART (Ahmad et al., 2021), T5 (Raffel et al., 2020), and CodeT5 (Wang et al., 2021), leverage bidirectional encoding and autoregressive decoding to address a variety of SE challenges. Extensions such as CodeT5+ (Wang et al., 2023b), AlphaCode (Li et al., 2022a), and CoTexT (Phan et al., 2021) have demonstrated enhanced adaptability in tasks such as Bug Fixing and Code Summarization (Al-Kaswan et al., 2023; Gu et al., 2022; Mastropaolo et al., 2021b). Additionally, ATHENATEST, proposed by Tufano et al. (Tufano et al., 2020), adapts the BART framework (Lewis et al., 2019) to generate unit tests, showcasing the versatility of encoder-decoder architectures in supporting software quality assurance tasks.
Decoder-only models, such as PaLM (Chowdhery et al., 2023), GPT-3.5 (OpenAI, 2023), GPT-4 (Achiam et al., 2023), and LLaMA (Touvron et al., 2023), represent a further evolution in LLM design, offering highly effective generative capabilities. For instance, Codex, with 12 billion parameters, has demonstrated the ability to solve 72.31% of complex Python programming problems and serves as the backbone for tools such as GitHub Copilot (Chen et al., 2021; Minaee et al., 2024). GPT-4 continues to set benchmarks across multiple SE tasks, including Code Synthesis, Debugging, and Reasoning (Bubeck et al., 2023). Evaluations by Arakelyan et al. (Arakelyan et al., 2023) on Codex and CodeT5 have further validated their effectiveness in Code Generation and Code Summarization, even under distribution shifts. Among the specialized decoder-only models, CodeLlama is a prominent family of open-source LLMs derived from LLaMA 2 (Touvron et al., 2023). It has been further trained on a corpus of 500 billion tokens, including both code and natural language, making it highly effective for various code-related tasks (Zan et al., 2024; Chen et al., 2021; Xia et al., 2023a; Virk et al., 2024). Likewise, DeepSeek-Coder (DeepSeek-AI et al., 2024) has emerged as a strong competitor, featuring models ranging from 1B to 33B parameters that have been pre-trained on 2 trillion tokens, including extensive code-related data. Remarkably, DeepSeek-Coder has been shown to outperform larger models like GPT-3.5 (Brown, 2020), with its 6.7B parameter model even surpassing the 33B variant of CodeLlama in certain benchmarks.
The effectiveness of these models is further augmented by continuous improvements in both architecture and task-specific tuning methods. Innovations such as infilling objectives (Li et al., 2023a) and instruction tuning (Xu et al., 2024) have been integrated into models like StarCoder and CodeLlama, improving their performance on tasks such as Code Completion and Code Editing (Roziere et al., 2023). Additionally, LCMs have been effectively applied to domain-specific and low-resource programming languages, including Verilog (Liu et al., 2023a) and Haskell (van Dam et al., 2024), often utilizing transfer learning from high-resource languages like Python and Java to overcome data scarcity issues (Zhang et al., 2023c). Beyond their conventional applications in programming tasks, LCMs have also been leveraged in emerging areas such as Autonomous Agent Systems for workflow automation (Hong et al., 2024) and in assessing the quality of AI-generated code (Wei et al., 2023a).
In conclusion, LCMs have consistently demonstrated state-of-the-art performance across a wide range of SE tasks, including Requirements Analysis (Marques et al., 2024), Code Comprehension (Izadi et al., 2024; Ding et al., 2023), Code Generation (Li et al., 2022a; Mastropaolo et al., 2024b), Code Translation (Pan et al., 2023), Code Summarization (Sun et al., 2024; Mastropaolo et al., 2024a, 2022), Automated Program Repair (Fan et al., 2023a; Jiang et al., 2023; Ribeiro, 2023; Xia et al., 2023b), Code Smell Detection and various Software Testing activities (Wang et al., 2024; Tufano et al., 2020; Wang et al., 2024). The continued evolution of models such as CodeLlama, DeepSeek-Coder, and CodeT5 underscores the transformative impact of domain-specific LCMs in automating and enhancing modern software engineering practices.
2.2. Fine-Tuning Code Models
The effectiveness of LCMs largely depends on their architectural scale, which often includes millions to billions of parameters. These models undergo pretraining on extensive source code corpora before being further adapted to downstream tasks through fine-tuning. Fine-tuning strategies can be broadly categorized as follows:
Full Fine-Tuning (FFT): This method initializes a Pretrained Language Model (PLM) with pretrained weights and refines it using task-specific data via backpropagation and gradient descent (Tayaranian et al., 2022). In FFT, all model parameters, including the pretrained weights, are updated to minimize a task-specific loss function that quantifies the deviation between model predictions and ground truth. However, full fine-tuning is computationally intensive, requiring substantial computational resources and large labeled datasets. The challenge becomes even more pronounced as PLMs scale to billions of parameters, making FFT impractical for many applications (Patil and Gudivada, 2024).
Parameter-Efficient Fine-Tuning (PEFT): To address the resource-intensive nature of FFT, researchers have developed PEFT methods, which update only a small subset of model parameters while keeping the rest frozen (Houlsby et al., 2019). This substantially reduces both memory requirements and computational costs, making it a more feasible approach for adapting LLMs to specific tasks. Recent studies have demonstrated the effectiveness of PEFT in software engineering applications, including Code Generation (Weyssow et al., 2023), Code Review (Lu et al., 2023), Code Clone Detection (Ayupov and Chirkova, 2022), and Automated Program Repair (Li et al., 2024b). These studies highlight how PEFT significantly reduces training time and memory consumption while maintaining or even improving model performance in various software engineering tasks.
2.3. Parameter-Efficient Fine-Tuning (PEFT)
PEFT techniques can be categorized into Additive, Selective, Reparameterized, and Hybrid fine-tuning based on their underlying mechanisms.
2.3.1. Additive PEFT:
Additive PEFT methods introduce new trainable parameters while keeping the original model parameters frozen, making them the most widely explored category in PEFT research. Two major subcategories within this approach include Adapter-based methods and Soft Prompts.
Adapter-based approaches integrate small adapter layers within Transformer blocks to enable task-specific adaptation without modifying the core model parameters (Houlsby et al., 2019). An adapter layer typically comprises a down-projection matrix $\mathbf{W}_{\text{down}}\in\mathbb{R}^{r\times d}$, followed by a non-linear activation function $\sigma(\cdot)$, and an up-projection matrix $\mathbf{W}_{\text{up}}\in\mathbb{R}^{d\times r}$. Here, $d$ denotes the dimensionality of the hidden layer, while $r$ is the bottleneck dimension, a tunable hyperparameter that controls the adapter’s capacity. Given an input $h_{\text{in}}$ to the adapter, the transformation performed by the adapter layer (including the residual connection) can be expressed as:
$\text{Adapter}(x)=\mathbf{W}_{\text{up}}\sigma(\mathbf{W}_{\text{down}}x)+x.$
This methodology has been widely adopted (Houlsby et al., 2019; Rücklé et al., 2020; Lin et al., 2020; Lei et al., 2023; He et al., 2021) and has led to multiple variations, such as modifications in adapter placement for instance SparseAdapter (He et al., 2022), Pruning techniques to enhance efficiency (He et al., 2021), and Reparameterization strategies aimed at reducing the number of trainable parameters (Karimi Mahabadi et al., 2021).
Two prominent types of adapters are the Language Adapter (L-Adapter) and the Task Adapter (T-Adapter) (Goel et al., 2022). The L-Adapter is trained using a masked language modeling objective on unlabeled data from a specific target language, thereby enabling the PLMs to learn language-specific characteristics. In contrast, the T-Adapter is optimized on labeled data to perform a specific downstream task, facilitating task-specific adaptation. This modular design enables PLMs to be effectively adapted to previously unseen languages, even those not originally covered during the model’s pretraining phase.
Named Entity Recognition (NER) (Saberi et al., 2024) Adapters are lightweight modules integrated within the Transformer architecture. Their primary objective is to capture and encode type information derived from the Abstract Syntax Tree (AST) of source code. By incorporating structural and semantic cues from the AST, these adapters enhance the model’s ability to recognize and classify named entities within code, contributing to more accurate and context-aware representations.
AdapterFusion (Pfeiffer et al., 2020) is a framework that integrates multiple task-specific adapters into a unified adapter module, facilitating effective knowledge transfer across related tasks without altering the parameters of the original pretrained model. This approach enables efficient task composition, allowing pretrained language models to be reused across a variety of downstream tasks.
MAM Adapter (He et al., 2021) explores the intrinsic similarities among three additive Parameter-Efficient Fine-Tuning (PEFT) methods: adapters, prefix-tuning, and LoRA. This insight leads to the development of three adapter variants. The first variant, Parallel Adapter, places adapter layers alongside specific components of the model, such as the self-attention (SA) or feedforward network (FFN) layers, instead of inserting them sequentially after these components. The second, Multi-head Parallel Adapter, divides the parallel adapter into multiple heads, with each head independently modifying the output of a corresponding attention head in the SA module. The third, Scaled Parallel Adapter, introduces a scaling factor following the parallel adapter layer, aligning with the design principles of LoRA.
Soft Prompt Tuning, another prominent additive PEFT method, introduces trainable continuous vectors, referred to as soft prompts, into either the input representation or the hidden states of the model. This can be represented as follows:
$X_{plq}=\left(s_{plq_{1}},\dots,s_{plq_{N_{S}}},x_{plq_{1}},\dots,x_{plq_{N_{X}}}\right)$
where $X_{plq}$ is the sequence of input tokens for layer $l$, including soft prompt tokens $s_{plq_{i}}$ followed by the original input tokens $x_{plq_{i}}$. $N_{S}$ is the number of soft prompt tokens, and $N_{X}$ is the number of original input tokens. Unlike manually crafted hard prompts—static natural language instructions or templates written by humans—soft prompts are automatically optimized through task-specific training data, allowing the model to learn an optimal prompt configuration. Variants of this approach include Prefix Tuning, which pre-pends soft prompts to the hidden states of the multi-head attention mechanism (Li and Liang, 2021), and Prompt Tuning, which appends learnable prompt tokens to the input embeddings while ensuring that only these prompt-related parameters are updated during training (Lester et al., 2021). By keeping the core model parameters frozen, these techniques offer computationally efficient fine-tuning while maintaining strong task adaptability.
P-Tuning (Liu et al., 2024d) introduces a prompt-based fine-tuning strategy by inserting a sequence of continuous soft prompts, denoted as $P_{1}],\cdots,[P_{i}],[P_{i+1}],\cdots,[P_{l}]$, into the input of a pretrained language model. Unlike traditional Prompt Tuning, P-Tuning constructs a learnable template by concatenating these prompts and mapping them into the embedding space as follows:
$\{h_{1},\cdots,h_{i},e(x),h_{i+1},\cdots,h_{l},e(x)\},$
where $e(x)$ represents the output from the pretrained embedding layer and $\{h_{1},\cdots,h_{l}\}$ are the trainable continuous prompt representations. The training objective is to optimize these soft prompt embeddings while keeping the parameters of the PLM fixed. This approach significantly reduces the number of trainable parameters, making it particularly effective in few-shot learning scenarios.
Pass-Tuning (Chen et al., 2023b) is a structure-aware, parameter-efficient fine-tuning strategy specifically designed for code representation learning. This approach introduces a modular, plug-and-play Graph Neural Network (GNN) component, which functions as a tunable prefix and is trained to capture structural information from the Abstract Syntax Tree of source code. By explicitly modeling the syntactic and hierarchical relationships inherent in the AST, Pass-Tuning facilitates a more nuanced and semantically rich understanding of code structure. This enables the underlying language model to better encode structural dependencies, thereby enhancing the quality of the learned code representations while maintaining efficiency in parameter usage.
[…]
AdaMix (Wang et al., 2022a), AdapterDrop (Rücklé et al., 2020), Ladder-Side Tuning (LST) (Sung et al., 2022), and Intrinsic Prompt Tuning (IPT) (Qin et al., 2021) are other popular additive PEFT methods.
2.3.2. Selective PEFT:
Selective PEFT optimizes a subset of existing parameters rather than introducing additional components, reducing computational overhead while maintaining adaptability. Consider a model with parameters $\theta=\{\theta_{1},\theta_{2},\ldots,\theta_{n}\}$, where each $\theta_{i}$ denotes an individual model parameter, and $n$ is the total number of parameters. Selective PEFT operates by applying a binary mask $M=\{m_{1},m_{2},\ldots,m_{n}\}$ over these parameters, where each $m_{i}\in\{0,1\}$ indicates whether parameter $\theta_{i}$ is selected ($m_{i}=1$) or excluded ($m_{i}=0$) from fine-tuning. The updated parameter set $\theta^{\prime}$ after fine-tuning can be expressed as:
$\theta_{i}^{\prime}=\theta_{i}-\eta\cdot m_{i}\cdot\frac{\partial L}{\partial\theta_{i}},$
where $\eta$ is the learning rate, and $\frac{\partial L}{\partial\theta_{i}}$ represents the gradient of the loss function $L$ with respect to $\theta_{i}$. Under this formulation, only the parameters for which $m_{i}=1$ are updated during backpropagation, while the remaining parameters remain fixed.
It can be classified into three approaches: Bias update methods, such as BitFit (Bias-term Fine-tuning) (Zaken et al., 2021), which fine-tunes only bias terms and task-specific layers while keeping most parameters frozen; pretrained weight masking, such as FISH Mask (Fisher-Induced Sparse Hanging) (Sung et al., 2021), which applies structured pruning to mask pretrained weights; and delta weight masking (Ansell et al., 2021; Xia et al., 2022; Guo et al., 2020), which prunes delta weights using optimization techniques like LT-SFT (Lottery Ticket Sparse Fine-Tuning) (Ansell et al., 2021).
For Instance, Telly-K (Shi et al., 2023) method uses a layer-freezing strategy where the bottom $K$ layers of the model are frozen— i.e., their parameters remain unchanged—while only the top (LK)(L-K) layers are updated during training, where $L$ denotes the total number of transformer layers. For example, in the Telly-6 configuration, layers 0 through $5$ are fixed, and fine-tuning is applied exclusively to layers $6$ through $11$.
2.3.3. Reparameterized PEFT:
Reparameterized PEFT methods employ low-rank transformations to reduce the number of trainable parameters while preserving the model’s ability to operate on high-dimensional matrices, such as pretrained weights. This category includes intrinsic dimension-based approaches, such as Intrinsic SAID (Aghajanyan et al., 2020), which was one of the earliest techniques exploring the intrinsic dimensionality of fine-tuning LLMs. However, the most widely adopted reparameterization technique is LoRA (Low-Rank Adaptation) (Hu et al., 2022). LoRA optimizes fine-tuning by introducing two trainable low-rank matrices—a down-projection matrix and an up-projection matrix—which are applied in parallel to update model weights efficiently. Given a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, LoRA (Low-Rank Adaptation) introduces two additional trainable matrices: a down-projection matrix $\mathbf{W}_{\text{down}}\in\mathbb{R}^{r\times k}$ and an up-projection matrix $\mathbf{W}_{\text{up}}\in\mathbb{R}^{d\times r}$, where the rank $r\ll\min(d,k)$. These matrices operate in parallel to $\mathbf{W}_{0}$, forming a low-rank approximation of the weight update. Letting $h_{\text{in}}$ represent the input, the standard output is computed as
$h_{\text{out}}=\mathbf{W}_{0}h_{\text{in}}.$
LoRA modifies this computation by injecting a task-specific incremental update $\Delta\mathbf{W}$, resulting in:
$h_{\text{out}}=\mathbf{W}_{0}h_{\text{in}}+\alpha\mathbf{W}_{\text{up}}\mathbf{W}_{\text{down}}h_{\text{in}},$
where $\alpha$ is a scaling factor that controls the magnitude of the adaptation. During initialization, $\mathbf{W}_{\text{down}}$ is sampled from a Gaussian distribution, while $\mathbf{W}_{\text{up}}$ is initialized to zero, ensuring that $\Delta\mathbf{W}=0$ at the start of training. This design makes LoRA straightforward to implement, with demonstrated scalability on models with up to 175 billion parameters. Once fine-tuning is completed, the learned LoRA weights can be merged with the pre-trained weights $\mathbf{W}_{0}$, allowing the model to retain its efficiency and avoid additional inference overhead.
Several LoRA-based extensions have been proposed, including KronA (Kronecker Adapter) (Edalati et al., 2022), DyLoRA (Valipour et al., 2022) , GLoRA (Chavan et al., 2023), and AdaLoRA (Zhang et al., 2023a), each introducing variations to enhance efficiency and task adaptability.
AdaLoRA (Adaptive Low-Rank Adaptation) (Zhang et al., 2023a) extends the standard LoRA framework by introducing a mechanism for dynamically adjusting the rank of adaptation matrices, thereby enabling efficient control over parameter allocation budgets. In this approach, the incremental weight update $\Delta W$ is reparameterized using Singular Value Decomposition (SVD) as $\Delta W=P\Lambda Q$, where $P$ and $Q$ are orthogonal matrices, and $\Lambda$ is a diagonal matrix containing the singular values $\{\sigma_{1},\sigma_{2},\ldots,\sigma_{r}\}$. The variable $r$ denotes the rank of $\Lambda$. During training, the matrices $P$ and $Q$ are initialized with values sampled from a Gaussian distribution and regularized to preserve orthogonality, while $\Lambda$ is initialized to zero. The rank of $\Lambda$ is iteratively pruned by truncating its smaller singular values, allowing the model to adaptively control the complexity of $\Delta W$.
FF-LoRA (He et al., 2021) is a variant of the LoRA framework that applies the low-rank update mechanism specifically to the Feed-Forward (FF) layers of Transformer-based architectures. While the original LoRA approach introduces a learnable low-rank decomposition $\Delta W=AB$, added to the model weights post-finetuning, its application was initially restricted to the Query and Value projections within the Multi-Head Attention mechanism. FF-LoRA extends this methodology by applying the low-rank update to the FF layers, enabling parameter-efficient adaptation in a different architectural component.
A notable extension of LoRA, QLoRA (Quantized LoRA) (Dettmers et al., 2023), integrates quantization techniques to further reduce computational demands. QLoRA fine-tunes transformer models by applying 4-bit NormalFloat (NF4) quantization and double quantization processing, significantly improving memory efficiency while maintaining fine-tuning effectiveness.
2.3.4. Hybrid/Others Technique:
Hybrid fine-tuning combines multiple PEFT techniques, such as Adapters, Prefix Tuning, and LoRA, to maximize efficiency and adaptability. Notable approaches include MAM Adapter (He et al., 2021), Compacter (Karimi Mahabadi et al., 2021), and AutoPEFT (Zhou et al., 2024), which integrate diverse strategies for enhanced performance.
For Instance, MAM Adapter (Mix-And-Match) (He et al., 2021) explores the intrinsic similarities among three PEFT methods: adapters, Prefix Tuning, and LoRA. This insight leads to the development of three adapter variants. The first variant, Parallel Adapter, places adapter layers alongside specific components of the model, such as the self-attention (SA) or feedforward network (FFN) layers, instead of inserting them sequentially after these components. The second, Multi-head Parallel Adapter, divides the parallel adapter into multiple heads, with each head independently modifying the output of a corresponding attention head in the SA module. The third, Scaled Parallel Adapter, introduces a scaling factor following the parallel adapter layer, aligning with the design principles of LoRA.
Prior research has explored the application of PEFT techniques across various NLP tasks. Han et al. (Han et al., 2024) conducted a comprehensive literature survey on existing PEFT methods and their adaptation across different domains. Lialin et al. (Lialin et al., 2023) performed a comprehensive experimental evaluation of PEFT methods, benchmarking 14 techniques and their variants across five datasets and three model scales (0.7B, 3B, and 11B parameters). The study provides an in-depth analysis of the methods’ efficiency, focusing on GPU memory usage and throughput performance. Other studies (Lialin et al., 2023; Xu et al., 2023) have similarly examined PEFT frameworks from a broader perspective. However, a systematic review of PEFT techniques specifically within software engineering remains largely unexplored. This study aims to systematically collect, analyze, and synthesize research on PEFT applications in SE tasks, providing a structured understanding of their effectiveness and impact. Ultimately, this work provides actionable insights for both researchers and practitioners by identifying which PEFT strategies are most effective across various SE tasks, outlining current limitations, and highlighting promising avenues for future research on integrating efficient fine-tuning techniques into software engineering workflows.
3. Methodology: Study Selection and Analysis
In this section, we outline the structure of our systematic literature review, following the well-established guidelines proposed by Kitchenham et al. (Kitchenham et al., 2009).
Our review is driven by three key Research Questions (RQs), which shape the direction of our investigation and ensure a focused analysis of primary studies related to PEFT techniques in software engineering. Specifically, our SLR is guided by the following RQs:
  • RQ1: Which software engineering tasks have been explored using PEFT methods?
    This research question investigates the full spectrum of software engineering applications where PEFT techniques have been applied, providing an overview of their adoption in the field. In particular, our focus is on highlighting differences (if any) that may emerge across various types of SE tasks, such as generative tasks (e.g., code generation, summarization) versus non-generative tasks (e.g., code classification, clone detection). By examining how PEFT strategies are utilized in these distinct contexts, we aim to uncover patterns of usage, task-specific adaptations, and potential gaps in the current literature that warrant further exploration.
  • RQ2: Which deep learning models and architectures are commonly optimized with PEFT methods?
    This research question aims to identify the types of base models and architectural families that are commonly fine-tuned using PEFT techniques. Additionally, it explores how different PEFT methods are instantiated across various model types (e.g., encoder-only, decoder-only, encoder-decoder) to support software engineering tasks. The goal is to understand which architectural configurations are most “compatible” with PEFT optimization methods.
  • RQ3: How do PEFT methods impact task performance and training efficiency in software engineering tasks when compared against full fine-tuning?
    This research question investigates the impact of the best-performing PEFT method(s) on both task performance (e.g., BLEU, Accuracy) and efficiency (e.g., training time, memory usage) across generative and non-generative software engineering tasks. To ensure a fair comparison, we consider only those studies that present a one-to-one evaluation between PEFT-optimized models and their fully fine-tuned counterparts. This approach enables us to draw meaningful insights into the practical applicability of PEFT techniques in real-world SE settings. For each study, we record whether the selected PEFT method leads to improvements, degradations, or neutral effects in terms of both performance and efficiency.
Through our examination of these research questions, we aim to map the landscape of Parameter-Efficient Fine-Tuning in software engineering, highlighting both its capabilities and limitations. Our review methodology consists of the following steps:
  1. (1)
    Identifying and Selecting Primary Studies
    1. (a)
      Executing search queries across four databases using a defined publication year range (2019–2025).
    2. (b)
      Manually screening results for relevance to PEFT methods and SE tasks.
  2. (2)
    Study Selection and Filtering Process
    1. (a)
      Exclusion Criteria and Alignment Analysis: Applied exclusion criteria and conducted manual review to finalize 25 studies, resolving disagreements through consensus.
    2. (b)
      Snowballing and Manual Addition of Studies: Identified and included two additional studies through backward snowballing, resulting in a final set of 28 papers.
  3. (3)
    Data Extraction and Synthesis
    1. (a)
      Data Extraction: Conducted a high-level extraction of relevant study attributes aligned with our research questions.
    2. (b)
      Synthesis Approach: Categorized SE tasks as generative or non-generative and analyzed trends in PEFT application across task types and deep learning models, as well as architectures.
Figure 1. Paper Selection Process for the Systematic Literature Review on PEFT Techniques in Software Engineering Tasks. The pipeline consists of five stages: applying search strings to databases— A, exclusion criteria— B, manual analysis— C, snowballing— D, and final selection— E.
Fig. 1 illustrates our paper selection workflow, outlining the progression from the initial search to the final set of papers included in our SLR. The detailed methodology for each step in this process is described in Section 3.2 (Study Selection and Filtering Process), including our search strategy, venue selection, and the application of inclusion and exclusion criteria such as snowballing.
3.1. Identifying and Selecting Primary Studies
To properly address the RQs underpinning our investigation, we look for primary studies by defining a time frame, ranging from 2019 to 2025, which aligns with the rise and evolution of PEFT techniques as a widely adopted method for optimizing large models.
Next, we began our search by querying four major electronic databases IEEE Xplore, ACM Digital Library, Springer Link, and Google Scholar (via Publish or Perish software) using a tailored search string designed to capture studies involving PEFT in SE contexts. This initial step–depicted in Fig. 1- A, resulted in 1,146 forming the basis of our analysis [^2] All the retrieved studies are available for consultation in our online appendix (rep, 2025) studies forming the basis of our analysis..
After retrieving the initial set of studies, we applied a two-stage filtering process. First, we conducted a relevance screening to ensure that the papers explicitly addressed PEFT methods and their application to software engineering tasks. This screening was performed by the first author, who manually reviewed the initial pool of 1,146 studies. In cases of uncertainty regarding the relevance of a particular approach to the software engineering domain, the first author consulted with the second author. If disagreement persisted, the conflict was solved by including the third author of the investigation. This allowed us to reach a consensus. Next, we applied venue-based filtering to ensure the quality and impact of the selected research. To this end, we focused on top-tier conferences and journals across software engineering, programming languages, machine learning, and natural language processing–Fig. 1- B. Our venue selection was informed by CSRankings [^3] https://csrankings.org, a metrics-driven and widely accepted ranking of top computer science venues. From this source, we selected high-impact venues including:
  • Artificial Intelligence: AAAI, IJCAI
  • Machine Learning: ICLR, ICML, NeurIPS, KDD
  • Natural Language Processing: ACL, EMNLP, NAACL
  • Programming Languages: PLDI, POPL
  • Software Engineering: FSE, ICSE, ASE, ISSTA
In addition to the venues extracted from CSRanking, we also included several well-regarded software engineering conferences and journals that are not listed in CSRankings but are recognized in the community for their relevance and peer-reviewed contributions. These include journals such as the ACM Transactions on Software Engineering and Methodology (TOSEM), IEEE Transactions on Software Engineering (TSE), and Empirical Software Engineering (EMSE), and conferences such as the Mining Software Repositories (MSR), International Conference on Software Analysis, Evolution and Reengineering (SANER), International Conference on Program Comprehension (ICPC), the international conference on AI Foundation Models and Software Engineering (FORGE).
As shown in Fig. 2, our final collection of papers spans 16 venues, with the ASE, ICSE, and TOSEM contributing the largest number of relevant studies. While this approach may exclude some potentially relevant papers from less prominent venues, it ensures our review emphasizes high-quality, peer-reviewed research from both top-ranked and widely recognized publication forums.
To further ensure rigor, we deliberately excluded preprint repositories (e.g., arXiv) from our primary database search. Although such sources may contain cutting-edge findings, the lack of formal peer review could introduce validity threats. Nonetheless, let us anticipate that through our snowballing process, we considered making an exception for two highly cited arXiv papers (Ayupov and Chirkova, 2022; Goel et al., 2022) that had substantial influence in the peer-reviewed literature, demonstrating their relevance and impact. In doing so, we adopted an approach grounded in sound principles and informed by domain expertise—allowing us to holistically assess the quality of each study. This includes evaluating its scholarly impact, such as citation count, and its influence within the relevant research community.
With a general overview of the approach we followed to implement the SLR, let us now introduce the fundamental details underlying the retrieval phase–Fig. 1A.
We first design the search string based on key terms extracted from our RQs, ensuring comprehensive coverage of PEFT methods (e.g., LoRA, Adapters, Prefix Tuning) and their applications in SE tasks. To construct the SE task list used in our query, we additionally followed the taxonomy proposed by Watson et al. (Watson et al., 2022) which provides a broad categorization of SE tasks studied in the context of DL4SE—an emerging area that applies deep learning techniques to solve a variety of software engineering problems.
To guide our retrieval process, we explicitly incorporated all 23 SE task names from Watson’s taxonomy into our search string. This design choice ensured broad task coverage and enabled the retrieval of papers relevant to both generative and non-generative SE tasks. Nevertheless, as discussed in Section 3.2 and Section 3.3.1, several SE tasks from Watson’s taxonomy—such as Code Smells, Software Security, and Developer Forum Analysis—were not represented in our collection of papers. In other words, not all the software engineering tasks have been explored in the context of PEFT. A result that, although was expected, indicates that there is still untapped potential to explore such optimization methods in broader SE domains. We present the results of our SLR in Section 4.
We incorporated all SE task names reported in Table 1, which represent the tasks identified through our complete study selection, filtering, and snowballing process (Section 3.2).
For Google Scholar, we used Publish or Perish (PoP) software (Harzing, 2019). However, when using (PoP) software to retrieve Google Scholar papers, the search bar has a 256-character limit. Due to this constraint, we split our search string into seven separate queries, each covering different subsets of PEFT techniques and SE tasks. This ensured that our search remained exhaustive while adhering to PoP’s query limitations. A few concrete examples are provided below–while all the queries can be retrieved from the replication package of our study (rep, 2025).
  • (’Parameter Efficient Fine Tuning’ OR ’PEFT’ OR ’LoRA’ OR ’QLoRA’ OR ’Adapters’ OR ’Soft Prompts’ OR ’Prefix Tuning’ OR ’Prompt Tuning’) AND (’Software Engineering’ OR ’Automated Code Review’ OR ’Static Analysis’ OR ’Code Clone Detection’ OR ’Bug Prediction’ OR ’Commit Message Generation’)
  • (’Parameter Efficient Fine Tuning’ OR ’PEFT’ OR ’LoRA’ OR ’QLoRA’ OR ’Adapters’ OR ’Soft Prompts’ OR ’Prefix Tuning’ OR ’Prompt Tuning’) AND (’Code Comprehension’ OR ’Source Code Retrieval’ OR ’Traceability’ OR ’Bug-fixing Process’ OR ’Code Smells’)
  • (’Parameter Efficient Fine Tuning’ OR ’PEFT’ OR ’LoRA’ OR ’QLoRA’ OR ’Adapters’ OR ’Soft Prompts’ OR ’Prefix Tuning’ OR ’Prompt Tuning’) AND (’Software Testing’ OR ’Non-code-related Software Artifacts’ OR ’Software Energy Metrics’ OR ’Program Synthesis’)
3.2. Study Selection and Filtering Process
Building upon the initial filtering steps described in Section 3.1, we now detail the exclusion and inclusion criteria used to finalize our set of primary papers. These criteria were applied to ensure that each study addressed both PEFT methods and SE tasks, and that it was published in top-tier venues within our defined time frame (2019-2025).
Figure 2. Venue Distribution of PEFT Techniques in Software Engineering.
3.2.1. Exclusion Criteria and Alignment Analysis
Following our inclusion criteria, we applied a set of exclusion rules to filter out studies that fell outside the scope of our SLR. Specifically, studies were excluded if they: (i) did not focus on SE-related tasks, (ii) were not published in one of the selected venues, or (iii) mentioned PEFT only as part of future work without presenting an actual implementation or evaluation. After applying these criteria to the initial 1,146 retrieved records, we narrowed the selection down to 26 papers that satisfied all requirements.
To ensure the reliability and quality of this selection, two independent reviewers (the first and the second author) conducted a detailed manual analysis of the 26 studies. Each reviewer systematically assessed the abstract, methodology, and contributions of the papers to verify whether they explicitly implemented, evaluated, or made a substantive contribution to PEFT techniques in the context of SE tasks. During this review, four disagreements emerged two related to inclusion and two to exclusion decisions. These cases were discussed collaboratively, and a consensus was reached to include all four disputed studies. Due to access limitations, one paper (Li et al., 2024a) was unavailable for analysis. Therefore, our investigation proceeded with the remaining 25 papers in the corpus–representing the outcome of Phase 3 of our SLR, also indicated with Fig. 1C.
3.2.2. Snowballing and Manual Addition of Studies
After finalizing the initial set of papers, we applied backward snowballing to identify additional relevant studies that may have been missed during the database search. This involved reviewing the references of selected peer-reviewed papers, which led to identifying two additional arXiv studies that met our inclusion criteria. This stage of the SLR is depicted in Fig. 1D.
We would like to remind the reader that, although these studies were not peer-reviewed at the time of inclusion, their presence in the reference lists of several reputable publications highlights their relevance and impact within the field. We therefore included the work by Goel et al. (Goel et al., 2022), cited by three related studies (Saberi and Fard, 2023; Saberi et al., 2024; Liu et al., 2023b), and the work by Ayupov et al. (Ayupov and Chirkova, 2022), cited by four studies (Fan et al., 2024; Chen et al., 2023b; Saberi et al., 2024; Liu et al., 2023b), both of which have been discussed or cited in multiple SE-focused PEFT papers (Fan et al., 2024; Chen et al., 2023b; Weyssow et al., 2023).
After incorporating these two additional studies, our final study set comprised 28 papers. The distribution of these papers across publication venues is shown in Fig. 2.
3.3. Data Extraction and Synthesis
Having finalized our corpus of 28 primary studies through rigorous search, filtering, and snowballing procedures (as described in Section 3.1 and Section 3.2), we proceeded to extract and organize data to answer our research questions. In the following subsections, we first describe our data extraction process and then outline the synthesis approach, including how we grouped software engineering tasks into generative and non-generative categories.
3.3.1. Synthesis Approach
As introduced in Section 1, we aim to provide evidence regarding the use of PEFT optimization for large code models in both generative and non-generative SE tasks. We define generative tasks as activities in which models produce or transform code artifacts or generate related textual outputs—such as natural language summaries of code. These tasks typically involve generating sequences of tokens, which represent fundamental outputs that large code models produce for both natural language and code-related applications. In contrast, non-generative tasks focus on analyzing existing code by interpreting, classifying, retrieving, or detecting specific properties. Examples include Defect Detection, Code Clone Detection, Code Search, and Method Name Consistency Check. This task classification serves as the foundation for our subsequent analysis, enabling us to map PEFT methods effectively to different types of software engineering tasks.
3.3.2. Data Extraction
To systematically address our three research questions, we followed a structured extraction protocol that enabled systematic collection of relevant attributes from each of the 28 selected primary studies.
RQ1 (Software Engineering Tasks): For each study, we identified the primary SE tasks being addressed, categorizing them as either generative (e.g., Code Generation, Code Summarization) or non-generative (e.g., Defect Detection, Code Search). Where available, we also recorded high-level task attributes, such as input/output formats and application contexts, to characterize how PEFT methods were applied across different SE tasks. This classification facilitates subsequent analyses for RQ2 and RQ3.
RQ2 (Model Architectures and PEFT Methods): We documented the base models used in each study, recording their architecture type (encoder-only, decoder-only, or encoder-decoder), model scale (e.g., up to 70B parameters), and specific Transformer variants (e.g., CodeT5-base, CodeLlama-13B). We also extracted information on PEFT methods employed in each SE task, including LoRA, QLoRA, Prefix Tuning, Adapters, Prompt Tuning, among others. Configuration-level details for commonly used PEFT methods were also collected, focusing on method-specific parameters and hyperparameters such as rank and the number of target layers, as well as general attributes, including the proportion of trainable parameters per task category.
RQ3 (Performance and Efficiency): We focused on comparing performance and computational efficiency between PEFT-optimized large code models and fully fine-tuned baseline models. For each study, we identified the best-performing PEFT methods (multiple methods if present) and recorded whether reported outcomes indicated improvements ( P E), equivalence ( P E), or declines ( P E) in performance and efficiency compared to the full fine-tuning baseline. Efficiency metrics such as reduced training time and memory consumption were noted where available. These findings were synthesized into a categorical summary, enabling task-wise comparisons of PEFT impacts on performance and efficiency.
Overall, our review encompasses 28 studies covering 19 distinct SE tasks, which we classify into twelve (12) generative and seven (7) non-generative task groups. These are further detailed in Table 1, where we report the task category in the leftmost column, followed by a full description of the software engineering task type–adopting the terminology established by Watson et al. (Watson et al., 2022). Finally, the third column reports the corresponding studies in which each task type was identified.
Table 1. Software Engineering Tasks Taxonomy into Generative and Non-Generative Categories with Corresponding Paper List.
Category SE Task Task Description Papers
Generative Automated Program Repair (APR) Automatically generates patches to fix buggy code without human intervention. (Li et al., 2024b; Liu et al., 2024c; Huang et al., 2025)
Code Generation (CG) Generates source code from natural language descriptions or formal specifications. (Wang et al., 2025b; Liu et al., 2024a; Fakih et al., 2024; Hajipour et al., 2024; Choi and Lee, 2023; Chen et al., 2023b; Shi et al., 2023; Ayupov and Chirkova, 2022; Yang et al., 2024a), (Wang et al., 2025a)
Code Refinement (CRF) Improves or restructures existing code while preserving its original functionality. (Peng et al., 2024; Hajipour et al., 2024; Chen et al., 2023b; Saberi and Fard, 2023)
Code Review Generation (CRG) Generates natural language review comments from code changes to provide immediate feedback during commits. (Fan et al., 2024)
Code Summarization (CSum) Generates concise natural language summaries describing code behavior or intent. (Choi and Lee, 2023; Xie et al., 2023; Saberi and Fard, 2023; Chen et al., 2023b; Saberi et al., 2024; Shi et al., 2023; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b, 2023a; Afrin et al., 2025; Liu et al., 2024a)
Code Translation (CTrans) Converts code from one programming language to another while preserving functionality. (Choi and Lee, 2023; Chen et al., 2023b; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b; Liu et al., 2024a), (Wang et al., 2025a)
Commit Message Generation (CMG) Generates commit messages automatically from code changes to summarize their intent and content. (Liu et al., 2024b; Fan et al., 2024)
Code Completion (CComp) Predicts and completes the next code token or line based on the surrounding context. (Shi et al., 2023; Liu et al., 2024a), (Wang et al., 2025a)
Just-in-time Comment Update (JITCU) Automatically updates code comments in response to changes in the associated code. (Fan et al., 2024)
Method Name Recommendation (MNR) Suggests meaningful method names based on code content and contextual information. (Zhu et al., 2023)
Protocol Buffer Transformation (PBT) Generates code functions that transform data structures between different format specifications. (Wang et al., 2025a)
Unit Test Generation (UTG) Automatically generates unit test cases from source code to verify its correctness. (Zlotchevski et al., 2022; Liu et al., 2024a)
Non-Generative Cloze Test (CT) A probing task used to assess a model’s understanding of code syntax without updating its parameters. (Saberi et al., 2024; Goel et al., 2022)
Code Clone Detection (CCD) Identifies syntactically or functionally similar code fragments to reduce redundancy and improve maintainability. (Chen et al., 2023b; Saberi et al., 2024; Shi et al., 2023; Liu et al., 2023b; Ayupov and Chirkova, 2022; Goel et al., 2022)
Code Review (CR) Assists the review process by recommending reviewers or identifying issues in code changes. (Yu et al., 2024)
Code Search (CS) Retrieves semantically relevant code snippets given a natural language query. (Shi et al., 2023; Wang et al., 2023c, a)
Defect Detection (DD) Predicts whether a code snippet contains bugs or security vulnerabilities. (Chen et al., 2023b; Liu et al., 2023b, 2024b; Wang et al., 2023c, 2022b)
Header File Prediction (HFP) Predicts the correct file paths for header dependencies given source file information. (Wang et al., 2025a)
Method Name Consistency Check (MNCC) Determines whether a method name aligns with its implementation to detect naming inconsistencies. (Zhu et al., 2023)
While Table 1 outlines the range of software engineering tasks covered in our SLR, grouped by output type (generative vs. non-generative), Fig. 3–as anticipated in Section 1 expands on this by presenting the taxonomy that visually captures how PEFT methods are applied across task categories, offering a structured lens through which to assess current practices. By mapping adaptation techniques to task types, it highlights usage trends, gaps in coverage, and opportunities for more targeted or diversified applications of PEFT in software engineering.
Figure 3. PEFT Taxonomy for Software Engineering Tasks.
4. Results
In this section, we present and analyze the findings of our investigation, structured according to the previously defined Research Questions. Each subsection corresponds to a specific RQ, offering a detailed examination of the results in relation to the objectives outlined in our study.
4.1. RQ1: Which software engineering tasks have been explored using PEFT methods?
This research question investigates and quantifies how PEFT methods are applied to enhance the efficiency, effectiveness, or both, of software engineering tasks. From our analysis of 28 primary studies, we identified 19 distinct SE tasks where PEFT-based techniques have been applied. These tasks have already been categorized into generative and non-generative types in Section 3.3.1. Generative tasks typically involve creating or manipulating code-related artifacts as output (e.g., source code, code documentation), often in an open-ended or token-by-token fashion. This includes tasks like code summarization, code generation, and commit message generation, even if the latter involves some form of templating–because the model must still synthesize content beyond merely selecting from fixed classes. On the other hand, non-generative tasks are those where the output space is bounded or categorical, such as classification, detection, or retrieval. For example, code classification–even though it produces an output–falls under the non-generative category because it selects from a finite set of labels and does not require constructing novel sequences of tokens. We further organize the 12 generative tasks based on their input-output modalities as follows: NL2Code tasks–which involve generating code from natural language, such as Code Generation and Unit Test Generation–appear in 12 studies; Code2NL tasks–which involve translating code into natural language, such as Code Summarization, Commit Message Generation, and Code Review Generation–appear in 16 studies; and Code2Code tasks–which involve code-to-code transformations or analysis, such as Code Translation, Code Refinement, Automated Program Repair, and others–are explored in 21 studies. While generative tasks are categorized based on their input-output modalities due to the diversity and structural implications of the output (i.e., text or code), non-generative tasks do not require such fine-grained modality breakdown. This is because the output across non-generative tasks is typically constrained to label spaces (i.e., binary, multi-class, regression values). To enhance clarity and better reflect the structural diversity of non-generative tasks, we classify them according to standard machine learning output paradigms (Zhang et al., 2023b; Peña and Herbold, 2025). Specifically, we group them into the categories as follows – Classification: Tasks where the model predicts a discrete class label, such as Code Clone Detection, Defect Detection and Code Classification; Regression: Tasks involving continuous-valued outputs; Ranking: Tasks that return ordered lists or scores, such as Code Search or Traceability Link Recovery.
We obtained 17 papers where more than one SE task was supported via PEFT methods (Hajipour et al., 2024; Choi and Lee, 2023; Zhu et al., 2023; Chen et al., 2023b; Saberi and Fard, 2023; Fan et al., 2024; Saberi et al., 2024; Shi et al., 2023; Liu et al., 2024b, 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b, 2023a; Goel et al., 2022; Liu et al., 2024a; Wang et al., 2025a)..
To ensure consistency in our classification, we clarify the following: when a paper applied a PEFT method to optimize a non-generative task that was used as part of a larger generative pipeline, we attributed the primary PEFT contribution to the non-generative task. This is because the PEFT technique was originally designed and evaluated for improving that specific component. The same principle applies in reverse: if a PEFT method was developed to enhance a generative task, even if it was later used within a broader classification or analysis framework, we still classified the contribution under the generative task.
This approach allows us to build a taxonomy that accurately reflects the original goal and scope of each PEFT application. Between 2022 and 2025, research on PEFT in software engineering has grown significantly, with an increasing number of papers addressing diverse tasks. This growing trend is depicted in Fig. 4 and highlights an early traction of Code Summarization (13 studies), and Code Clone Detection (6 studies), while Code Generation (10 studies) has only seen a notable rise in PEFT supports recently. Nonetheless, since our collection of relevant studies concluded on February 2025, we expect that additional papers-particularly in the areas of software engineering and broader applications-will be published by the end of the current year (2025).
Upon closer examination of generative tasks, we observe that Code Summarization accounted for $\sim$46.4% (13 studies) and Code Generation with 10 studies ($\sim$35.7%) are the most frequently explored tasks, reflecting strong interest in improving model efficiency for synthesis-based application. Prefix Tuning and Prompt Tuning are among the earliest and most extensively tested methods for adapting Transformer-based models to Code Summarization with minimal trainable parameters. Adapter-based techniques are also popular, as evidenced in (Chen et al., 2023b; Ayupov and Chirkova, 2022; Liu et al., 2023b; Wang et al., 2023a).
Furthermore, recent advancements in reparameterization-based methods (see Fig. 3)–such as QLoRA—have shown promising results. For instance, Afrin et al. (Afrin et al., 2025) employed QLoRA to fine-tune large-scale code models, including CodeLlama-7B and DeepSeek-Coder-33B, on the CodeXGLUE benchmark (Lu et al., 2021), targeting the task of code summarization. The proposed approach attained performance comparable to that of fully fine-tuned counterparts, while demonstrating substantial gains in memory efficiency.
Ayupov et al. (Ayupov and Chirkova, 2022) explores various PEFT methods like Adapter Tuning (AT), LoRA, FF-LoRA (He et al., 2021), and FF-LoRA + AT to fine-tune CodeT5-base and PLBART-base on the CodeSearchNet dataset (Husain et al., 2019), enabling parameter-efficient adaptation across diverse programming languages such as Python and Go–highlighting the interest from researchers of the field in exploring PEFT methods to cover SE-related activities that are developed to produce both type of information LCMs handle, i.e., code and technical natural language.
To this extent, we want to highlight the usage of Reparametrized methods and particularly, LoRA for optimizing models primed for the production of code tokens. For example, Liu et al. (Liu et al., 2024a) utilizes LoRA and QLoRA to fine-tune models such as QWen-14B and CodeLlama-Python-13B across benchmarks like HumanEval and MBPP, enabling efficient large-scale adaptation.
Not only have low-rank transformation methods been used for Code Generation, but additive approaches such as Prefix Tuning have also been explored. In this regard, Choi et al. (Choi and Lee, 2023) introduces an advanced Prefix Tuning method with input-dependent prompt templates and multi-word initialization to enhance CodeT5’s performance when generating Java code from natural language intents.
Moving forward, Code Translation also shows growing traction with 8 different papers exploring the extent to which PEFT techniques, including Additive, Reparametrized, as well as Hybrid, can serve the purpose of effective and efficient fine-tuning. For example, Liu et al. (Liu et al., 2023b) evaluates multiple PEFT methods including Adapter, LoRA, Prefix Tuning, Prompt Tuning, and Mix-And-Match (MAM) on CodeT5 and PLBART to translate code segments from Java to C#. In a different but related study, Wang et al. (Wang et al., 2023c) explore the capabilities of Prompt Tuning with varying sizes of CodeT5.
Within the domain of Code2Code tasks, Code Refinement has emerged as a prominent area of research, with 4 studies investigating the advantages and limitations of applying PEFT techniques to the challenge of fixing buggy code. Peng et al. (Peng et al., 2024), for example, examine the effectiveness of prompt-based tuning on a CodeT5-based model for bug-fixing tasks. Similarly, Saberi et al. (Saberi and Fard, 2023) employ both language-specific and task-specific adapters in combination with AdapterFusion to fine-tune CodeBERT, enabling it to generate meaningful fixes for buggy Java methods (Tufano et al., 2019). In parallel, Chen et al. (Chen et al., 2023b) explore a range of PEFT techniques—including Pass-Tuning, BitFit, Adapters, and P-Tuning—built on top of CodeT5 and PLBART to correct faulty code components (Tufano et al., 2019).
Similar findings are observed in the closely related area of Automated Program Repair, which includes three studies (Li et al., 2024b; Liu et al., 2024c; Huang et al., 2025) employing both Additive and Reparameterized PEFT methods. The Additive approaches include techniques such as (IA)3, P-Tuning, and Prefix-Tuning, while the Reparameterized category features methods like AdaLoRA and BaseLoRA. Notably, we identified that within each category, a specific PEFT variant—(IA)3 for Additive methods and AdaLoRA for Reparameterized ones—was exclusively applied to APR tasks, suggesting a targeted use of these techniques for this particular application.
Additionally, we noted that tasks such as Code Review and Code Review Generation have received increasing attention from researchers (Fan et al., 2024; Yu et al., 2024; Lu et al., 2023). Through our SLR, we identified two distinct formulations of code review activities. The first, Code Review Generation, is a generative task (as described in Table 1) that focuses on producing natural language feedback a reviewer would typically provide for a code change. The second, which we refer to as the Code Review, belongs to the non-generative family and involves identifying or analyzing aspects of the code review process—such as classifying review comments, linking reviews to commits, or detecting review coverage patterns. In this regard, Fan et al. (Fan et al., 2024) explored LoRA and Prefix Tuning on models such as InCoder and CodeGen, using a real-world diff-review dataset comprising over 138,000 instances. This highlights how PEFT methods can be effectively applied even in less frequently studied yet practically impactful SE tasks like Code Review Generation.
Other explored tasks include Just-in-time Comment Update, Method Name Recommendation, Code Completion, Protocol Buffer Transformation and Unit Test Generation. Across these tasks, LoRA, Prefix Tuning, and Adapters remain the most frequently used methods, often applied to large-scale models such as CodeLlama (Roziere et al., 2023), CodeT5 (Wang et al., 2021), and InCoder (Fried et al., 2022).
Similarly, in Commit Message Generation, Fan et al. (Fan et al., 2024) applies LoRA and Prefix Tuning to LLaMA and CodeGen models on the large-scale MCMD dataset by Tao et al. (Tao et al., 2022), which includes over 2 million commit-message pairs from top GitHub repositories.
When PEFT methods are applied to fine-tune large code models for method name recommendation, compelling evidence is provided by Zhu et al. (Zhu et al., 2023). In their study, Prompt Tuning is initialized using CodeT5 and enhanced by incorporating contextual code elements. The authors reframed the method name consistency check as a classification problem, leading to improved naming accuracy. This example illustrates how PEFT techniques can enable efficient and targeted adaptation, even for less frequently explored software engineering tasks.
Moving beyond generative tasks, PEFT techniques have also been explored for non-generative SE applications, with Code Clone Detection being the most prominent among them, appearing in 6 studies ($\sim$21.4%). Analyzing the distribution of non-generative tasks across machine learning paradigms reveals clear patterns. Classification tasks dominate, with 16 studies covering Cloze Test (2), Code Clone Detection (6), Code Review (1), Defect Detection (5), Method Name Consistency Check (1), and Header File Prediction (1). Ranking tasks appear in only 3 studies, all in Code Search applications, and no regression-based tasks were identified. Overall, approximately $\sim$84% of non-generative PEFT applications focus on discrete classification problems, split into 8 papers on binary classification ($\sim$66.7%), 2 on multi-class classification ($\sim$16.7%), and 2 addressing both ($\sim$16.7%). To this end, a relevant approach is Telly-K, proposed by Shi et al. (Shi et al., 2023), which selectively tunes the bottom layers of UniXcoder and GraphCodeBERT to detect Java code clones on the BigCloneBench dataset (Svajlenko and Roy, 2016) which is Defect Detection is another important non-generative task where PEFT methods have shown promise, appearing in 5 studies. This task focuses on identifying buggy code, often using C or Java datasets like Devign (Zhou et al., 2019) and JIT-Defects4J. For example, Liu et al. (Liu et al., 2024b) explores Adapter Tuning and LoRA on a diverse set of models—including CodeBERT, GraphCodeBERT, and PLBART using the JIT-Defects4J dataset with over 16K training instances to detect Just-in-time Software Defects. Similarly, Wang et al. (Wang et al., 2023c) applies Prompt Tuning to CodeBERT and CodeT5 variants on the Devign dataset, which includes over 27K C code snippets with nearly half labeled as defective. These efforts highlight how PEFT enables effective adaptation for real-world bug detection scenarios with limited parameter updates.
It is worth noting that Code Search appears in 3 studies, primarily evaluated on the CodeSearchNet dataset (Husain et al., 2019)–while for Cloze Test, which appears in 2 studies, we find applications of L-Adapter (Goel et al., 2022) trained on CodeSearchNet and CodeNet (Puri et al., 2021). Method Name Consistency check (1 study) is addressed by Zhu et al. (Zhu et al., 2023) using Prompt Tuning with CodeT5, where contextual prompts are crafted to convert the problem into a classification task, effectively enhancing consistency evaluation.
Figure 4. Trends of PEFT Techniques Applied to Software Engineering Tasks Over the Years. Generative tasks are marked with a star ($\star$), while non-generative tasks are marked with a diamond ($\diamond$). Note that 2025 is ongoing, and additional studies may still emerge throughout the year.
In summary, our analysis reveals a growing and diverse adoption of PEFT techniques across the landscape of software engineering tasks, encompassing both generative tasks (e.g., Code Summarization, Code Generation, Code Translation) and non-generative tasks (e.g., Defect Detection, Code Clone Detection, and Code Search). This widespread applicability highlights PEFT’s flexibility in adapting large code models with minimal parameter updates, making them more accessible for practical use in real-world software engineering scenarios where computational constraints are common.
{iconbox}
Key Findings (RQ1)[Uncaptioned image]
PEFT techniques have been applied to 19 diverse SE tasks, with a strong focus on generative tasks and growing interest in non-generative ones. While tasks like Code Summarization and Generation remain the most studied, several emerging tasks have also seen initial exploration. Common PEFT methods include Prompt Tuning, Adapter, LoRA, (IA)3, and QLoRA, applied across popular models such as CodeT5, CodeLlama, and CodeBERT. These findings demonstrate the adaptability and effectiveness of PEFT in software engineering with different model architectures and varying sizes. Task concentration patterns reveal that Code Summarization represents $\sim$46.4% of studies while Code Generation accounts for $\sim$35.7%, indicating clear research priorities in synthesis-based applications. Cross-task versatility analysis shows that $\sim$60.7% of studies apply PEFT to multiple SE tasks, demonstrating broader methodological flexibility than single-task approaches. The distribution across task categories shows non-generative tasks like Code Clone Detection representing $\sim$21.4% of studies, revealing systematic coverage of both generative and non-generative software engineering activities.
{iconbox}
Future Directions (RQ1)[Uncaptioned image]
Future research should expand PEFT applications to underexplored and emerging software engineering tasks that remain less represented in current literature. This includes areas such as testing, documentation, and code quality analysis, as well as forward-looking tasks like API Recommendation and Automated Refactoring. Another promising direction is cross-task adaptation, where a single PEFT configuration is designed to generalize across multiple activities. Additionally, investigating task-specific PEFT variants can offer insights into which methods work best for particular task types. Such efforts would help bridge research innovations with practical software development needs.
4.2. RQ2: Which deep learning models and architectures are commonly optimized with PEFT methods?
In RQ1, we examined the range of software engineering tasks that have been addressed using PEFT methods, highlighting trends over time. Building on that foundation, this section shifts focus to the deep learning models and particularly Transformers architectures (Vaswani et al., 2017) optimized with PEFT techniques across various software engineering tasks. In particular, we analyze the types of Transformer models that have been optimized using PEFT and their alignment with different development activities, focusing on whether certain architectures (e.g., encoder-only, decoder-only, encoder-decoder) are more commonly employed for PEFT-optimization in comparison to others. Furthermore, we examine how various PEFT methods (e.g., LoRA, QLoRA, Prefix Tuning, Adapters) are applied across different model architectures, and whether certain learning algorithms or optimization strategies consistently achieve better results. Ultimately, by analyzing these factors, we aim to identify patterns that illustrate how PEFT techniques are shaping the adaptation of deep learning models—particularly Transformer-based architectures—for software engineering applications.
4.2.1. Base Models and Architectures in PEFT Software Engineering Optimization.
Before detailing the results of this analysis, we remind the reader that one paper can employ more than one techniques, as a results of that, we found that encoder-decoder architectures are the most prevalent in PEFT-based software engineering optimization, appearing in 29 software engineering task instances across 12 studies (Hajipour et al., 2024; Choi and Lee, 2023; Zhu et al., 2023; Xie et al., 2023; Chen et al., 2023b; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b, 2023a; Peng et al., 2024; Zlotchevski et al., 2022). This is followed by encoder-only architectures with 22 applications across 10 studies (Hajipour et al., 2024; Saberi and Fard, 2023; Saberi et al., 2024; Shi et al., 2023; Liu et al., 2024b, 2023b; Wang et al., 2023c, 2022b, a; Goel et al., 2022), while we found that auto-regressive models (i.e., decoder-only architectures) are employed to automate 19 software engineering activities across 12 studies (Li et al., 2024b; Liu et al., 2024a; Fakih et al., 2024; Yang et al., 2024a; Hajipour et al., 2024; Fan et al., 2024; Yu et al., 2024; Liu et al., 2024c; Afrin et al., 2025; Huang et al., 2025; Wang et al., 2025b, a). These findings indicate a clear preference for encoder-decoder architectures among researchers, likely due to their effectiveness in handling both understanding and generation tasks common in SE applications. It is important to note that our goal in examining model architectures is not to suggest that architecture type prescriptively determines the choice of PEFT method. Instead, this analysis aims to document how current practices reveal consistent patterns in how PEFT techniques are applied across model families. By empirically mapping the distribution of PEFT methods over different architectures, we capture how practical considerations–such as model structure, parameter scale, and task orientation–correlate with the adoption of particular strategies.
When looking at the base models, our analysis reveals that CodeT5 is the most frequently adapted LCMs, appearing in a total of 19 times among all SE studies across 8 studies (Hajipour et al., 2024; Choi and Lee, 2023; Zhu et al., 2023; Chen et al., 2023b; Liu et al., 2024b, 2023b; Wang et al., 2023a; Peng et al., 2024).
For the encoder-only model GraphCodeBERT, we report 10 software engineering task instances across 5 studies (Hajipour et al., 2024; Saberi et al., 2024; Shi et al., 2023; Liu et al., 2024b, 2023b). These tasks primarily fall into the Code2Code and Code2NL categories.
Moving forward, CodeBERT–an encoder-only transformer model, has been used in 9 software engineering task instances across 6 studies (Saberi and Fard, 2023; Saberi et al., 2024; Liu et al., 2024b, 2023b; Wang et al., 2023c, 2022b), while PLBART appears in 9 SE task instances across 6 studies (Saberi and Fard, 2023; Saberi et al., 2024; Liu et al., 2024b, 2023b; Wang et al., 2023c, 2022b), and UniXcoder is applied to 8 tasks across 3 studies (Shi et al., 2023; Liu et al., 2024b; Wang et al., 2023a).
When examining specific model variants, the CodeLlama family shows particularly flexible adoption, with CodeLlama-7B appearing in 7 SE instances across 5 studies (Li et al., 2024b; Fakih et al., 2024; Fan et al., 2024; Liu et al., 2024c; Afrin et al., 2025), CodeLlama-13B in 6 SE instances across 3 studies (Li et al., 2024b; Fan et al., 2024; Wang et al., 2025a), and CodeLlama-34B in 5 studies (Liu et al., 2024a; Fakih et al., 2024; Afrin et al., 2025; Huang et al., 2025; Wang et al., 2025a). Crossing tasks and models, we find out that CodeLlama-7B is used across diverse tasks, including Automated Program Repair, Code Generation, Commit Message Generation, and Code Summarization. CodeLlama-13B extends this usage to Code Generation, Code Completion, Code Translation, and Unit Test Generation, Protocol Buffer Transformation, while CodeLlama-34B supports more resource intensive tasks such as Code Completion and Automated Program Repair. This shows that the CodeLlama models are effective across a wide range of tasks, including NL2Code tasks like Code Generation, Code2NL tasks like Code Summarization and Code Review Generation.
This distribution highlights researchers’ strategic selection of models that balance generation capability with efficiency, such as training time and memory usage, leading to a growing emphasis on larger, specialized models like CodeLlama. A full analysis of these trade-offs—considering both functional performance and computational efficiency-is provided in Section 4.3.
Several recent studies have explored a range of model architectures in their evaluations, including StarCoder-15B (Huang et al., 2025; Liu et al., 2024a), DeepSeek-Coder variants (Li et al., 2024b; Liu et al., 2024a; Afrin et al., 2025; Wang et al., 2025a), and QWen-14B (Liu et al., 2024a), reflecting interest in evaluating both general purpose language models and those specialized for code related tasks. The architecture distribution reveals clear task-specific patterns. Encoder-decoder models (CodeT5, PLBART) maintain their dominance in generation-heavy tasks, and thus, they are naturally applied in the context of generative applications. On the other hand, encoder-only models (CodeBERT, GraphCodeBERT) remain preferred for analysis-oriented tasks. Finally, decoder-only models (particularly CodeLlama variants) are demonstrating growing adoption across both Code Generation and Automated Program Repair tasks, as evidenced by multiple studies (Li et al., 2024b; Liu et al., 2024c; Wang et al., 2025b; Liu et al., 2024a; Fakih et al., 2024; Hajipour et al., 2024; Fan et al., 2024). Decoder-only models appeared exclusively in generative tasks. Specifically, they are used in NL2Code tasks such as Code Generation (5 studies), Code2NL tasks such as Code Summarization, Commit Message Generation, and Code Review Generation (each reported in 1 study), and Code2Code tasks such as Automated Program Repair (2 studies), Code Translation, Code Refinement, Protocol Buffer Transformation and Just-in-Time Comment Update (each reported in 1 study). The evolution of model selection shows a marked progression from earlier focus on established architectures to current exploration of larger, more specialized models. This shift is particularly evident in the rising adoption of 13B+ parameter models like CodeLlama variants (7B/13B/34B), DeepSeek-Coder-33 (Liu et al., 2024a; Afrin et al., 2025; Wang et al., 2025a), and LLaMA-2-70B (Liu et al., 2024a), enabled by advancing PEFT techniques that make such models more accessible even in resource-constrained environments.
4.2.2. PEFT Methods Across Software Engineering Tasks and Model Architectures
In this section, we explore how PEFT methods are distributed across various software engineering tasks, distinguishing them between generative and non-generative. Among the identified techniques, Base LoRA is the most dominant, appearing in 29 SE tasks. The Adapter family has been applied across 23 software engineering tasks, encompassing specialized variants such as L-Adapters and T-Adapters. Other frequently used techniques include Prompt Tuning (14 studies), Prefix Tuning (13), and P-Tuning (7), which leverage soft or continuous prompts to inject task-specific knowledge into frozen pre-trained models. Collectively, these methods demonstrate significant effectiveness in reducing training costs while maintaining or enhancing performance, affirming their value in scalable and efficient model adaptation across a diverse set of software engineering tasks.
For generative tasks, we observe distinct patterns in the application of PEFT methods. In Automated Program Repair, three studies adopted LoRA (Li et al., 2024b; Liu et al., 2024c; Huang et al., 2025), highlighting an interesting correlation task-method. Additionally, one study (Li et al., 2024b) explored other PEFT strategies such as P-Tuning and Prefix Tuning independently, while another study evaluated AdaLoRA and (IA)3 for APR (Huang et al., 2025). This reflects a diverse experimentation landscape where different lightweight fine-tuning techniques are applied across studies to assess their effectiveness.
Code Generation is addressed in 10 studies, making it one of the most frequently explored generative SE tasks in the context of PEFT. LoRA remains the dominant technique, appearing in 7 studies (Wang et al., 2025b; Liu et al., 2024a; Fakih et al., 2024; Hajipour et al., 2024; Ayupov and Chirkova, 2022; Yang et al., 2024a; Wang et al., 2025a), including combinations of FF-LoRA variants with Adapter Tuning, one study implemented Code Summarization, Code Clone Detection, Code Generation, and Code Translation (Ayupov and Chirkova, 2022).
Despite the wide employment of Reparametrized techniques, it is important to acknowledge the community’s commitment to exploring the full spectrum of PEFT methods—ranging from Additive to Hybrid approaches—for Code Generation tasks. This growing body of work reflects a clear interest in identifying which PEFT strategies most effectively adapt large models to one of the most central and computationally demanding software engineering activities.
A further application of SE automation revolves around the generation of unit tests–referred to as Unit Test Generation in our taxonomy Fig. 3. In particular, for such a task we have found two distinct studies, each employing distinct PEFT approaches. Liu et al. (Liu et al., 2024a) applied QLoRA within a multi-task fine-tuning framework (MFT-5Tasks), while Zlotchevski et al. (Zlotchevski et al., 2022) employed Prefix Tuning, wherein 200 trainable virtual tokens were prepended to a frozen BART model.
Looking at SE tasks for which code changes optimization is requested, we find a surprisingly widespread application of PEFT techniques to support Code Refinement activities (i.e., bug-fixing). This peculiar task, has been explored through a diverse range of PEFT techniques, which—consistent with findings from Code Generation—span the full spectrum of parameter-efficient optimization strategies. In this context, Nuo et al. (Chen et al., 2023b) introduced Pass-Tuning, a lightweight method that tunes less than 1% of parameters by leveraging U-AST (Upgraded Abstract Syntax Tree) token sequences within a graph-attentional model. Building on structure-aware prompting, Yun et al. (Peng et al., 2024) applied Prompt Tuning to CodeT5-base, crafting domain-specific prompts based on AST-derived fix templates. Saberi et al. (Saberi and Fard, 2023) proposed a fusion-based approach, combining a Java-specific L-Adapter with a token-type-aware NER-Adapter via AdapterFusion to improve bug-fixing effectiveness. Similarly, Hajipour et al. (Hajipour et al., 2024) investigated the use of LoRA as a parameter-efficient alternative for Code Refinement tasks. Traditional Prompt Tuning appears in one study (Peng et al., 2024), while a multi-perspective adaptation strategy—incorporating L-Adapter, NER-Adapter, and AdapterFusion—is explored in a different research (Saberi and Fard, 2023). Additionally, one study applies a comprehensive suite of PEFT methods—including Pass-Tuning, BitFit, Adapters, and P-Tuning—to enhance source code refinement (Chen et al., 2023b).
Still in the domain of Code2Code tasks, Code Translation is addressed in 8 studies (Choi and Lee, 2023; Chen et al., 2023b; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b; Liu et al., 2024a; Wang et al., 2025a), while Code Completion is explored in 3 studies (Shi et al., 2023; Liu et al., 2024a; Wang et al., 2025a). Notably, for the latter, Shi et al. (Shi et al., 2023) introduce a customized PEFT technique called Telly-K, which modifies the training procedure by selectively freezing the lower layers of the model. This approach is designed to preserve foundational representations while reducing computational overhead during fine-tuning. Beyond these, several other PEFT techniques have been applied to Code2Code tasks, where code manipulation is central. Prefix Tuning is adopted in 2 studies (Choi and Lee, 2023; Liu et al., 2023b), while Prompt Tuning is utilized in 3 studies (Liu et al., 2023b; Wang et al., 2023c, 2022b).
Reparametrized methods such as LoRA and its variants—including FF-LoRA and FF-LoRA + AT—are employed in 2 further studies (Liu et al., 2023b; Ayupov and Chirkova, 2022). Additionally, Adapter-based strategies such as MAM, as well as lightweight approaches like Pass-Tuning, BitFit, and P-Tuning, are explored individually across different works (Chen et al., 2023b; Liu et al., 2023b; Ayupov and Chirkova, 2022).
In light of this analysis, a noteworthy trend emerges: the wide range of PEFT methods outlined above reflects the community’s active exploration of diverse strategies to determine the most effective parameter-efficient techniques for code transformation tasks. This ongoing experimentation highlights both the complexity of these tasks and the growing interest in tailoring fine-tuning approaches to their specific requirements.
Turning to Code2NL tasks, Code Review Generation is investigated in 1 study (Fan et al., 2024), which independently applies both LoRA and Prefix Tuning for efficient adaptation. Code Summarization, however, emerges as the most frequently studied Code2NL task, appearing in 13 studies. A broad set of PEFT techniques has been explored in this context. Among additive methods, Prefix Tuning and Adapter-based approaches—including L-Adapters, T-Adapters, and AdapterFusion—are used in 3 (Choi and Lee, 2023; Xie et al., 2023; Liu et al., 2023b) and 6 studies (Saberi and Fard, 2023; Chen et al., 2023b; Saberi et al., 2024; Liu et al., 2023b; Ayupov and Chirkova, 2022; Wang et al., 2023a), respectively. Prompt Tuning is featured in 3 studies (Liu et al., 2023b; Wang et al., 2023c, 2022b), while LoRA and its variants (FF-LoRA and FF-LoRA + AT) are applied in 2 studies (Liu et al., 2023b; Ayupov and Chirkova, 2022). Additional techniques such as Pass-Tuning, BitFit, P-Tuning, and MAM are also explored.
Notably, QLoRA is also employed in two Code2NL studies (Liu et al., 2024a; Afrin et al., 2025). In particular, Afrin et al. (Afrin et al., 2025) successfully applied QLoRA to fine-tune large language models with up to 4 billion parameters for the task of Code Summarization, demonstrating the feasibility of efficient adaptation at varying scales.
Commit Message Generation follows with 2 studies–both by relying on the application of LoRA (Fan et al., 2024; Liu et al., 2024b). However, while study combines LoRA with Prefix Tuning (Fan et al., 2024), the other integrates it with Adapter Tuning (Liu et al., 2024b).
Non-generative SE tasks typically adopt more targeted adaptation strategies, leveraging structured PEFT methods like Adapter and Prompt Tuning to enhance code understanding and classification. Cloze Test for example, is covered by 2 studies (Saberi et al., 2024; Goel et al., 2022), both employing Adapter-based methods.
Code Clone Detection has been explored across 6 studies, using a variety of PEFT methods, particularly, Adapter-based strategies. Study by Goel et al. (Goel et al., 2022) employs L-Adapter and T-Adapter modules trained on datasets such as Big Clone Bench (BCB), POJ-104, and SCD-88 to perform Code Clone Detection. Other studies (Chen et al., 2023b; Liu et al., 2023b; Ayupov and Chirkova, 2022) adopt modular PEFT strategies including Adapter, P-Tuning, Pass-Tuning, and BitFit.
Code Search is addressed in 3 studies, each applying distinct PEFT methods. A work by Wang et al. (Wang et al., 2023c), Prompt Tuning was applied using CodeBERT in a two-tower architecture, where hard prompts were designed separately for code and natural language inputs to generate embedding vectors for cosine similarity based retrieval. A study by Wang et al. (Wang et al., 2023a) use Adapter with UniXcoder, by configuring lightweight 5M-parameter setup that outperformed fully fine-tuned baselines such as CodeBERT and PLBART. Another study (Shi et al., 2023) explored PEFT methods alongside traditional architectures but is excluded here from detailed discussion, as its method (Telly-K) was previously covered. These works highlight the applicability of Prompt and Adapter-based strategies for retrieval tasks without adjusting the whole set of parameters describing the underlying model.
In the domain of non-generative tasks, Defect Detection has been explored in 5 studies, showcasing a diverse range of PEFT techniques aimed at identifying faulty or vulnerable code segments. Among these, additive methods appear to be the preferred choice, with Prompt Tuning adopted in 3 studies (Wang et al., 2023c, 2022b; Liu et al., 2023b), underscoring its effectiveness for classification-oriented tasks.
It is also worth highlighting that several studies implement multiple PEFT techniques within the same experimental setup—including Adapters, LoRA, and others (Liu et al., 2024b, 2023b)—although these are typically applied independently rather than in an integrated manner. Notably, Chen et al. (Chen et al., 2023b) evaluate a recurring combination of methods, including Pass-Tuning, BitFit, Adapter, and P-Tuning, which have also been observed in other software engineering tasks, suggesting their broad applicability across different adaptation scenarios.
Overall, while non-generative tasks employ a targeted subset of PEFT techniques, generative tasks display a broader range of strategies—with particularly high adoption of LoRA and prompt-based methods. This trend reflects the growing need for scalable, low-resource generative models in software engineering, supporting the increasing integration of generative AI across a wide array of development activities.
Table 2 summarizes the distribution of PEFT methods across generative and non-generative software engineering tasks, highlighting which specific papers employ each method. Software Engineering task abbreviations are defined in Table 1.
Table 2. PEFT Methods and Sub-Variants Across Software Engineering Tasks (Generative vs. Non-Generative). Note: Multiple SE tasks followed by a single citation indicate that all tasks were addressed in the same paper.
Main Method Sub-Method Generative Tasks Non-Generative Tasks
LoRA-based Base LoRA APR(Li et al., 2024b; Liu et al., 2024c; Huang et al., 2025), CComp(Liu et al., 2024a) (Wang et al., 2025a), CG(Fakih et al., 2024; Hajipour et al., 2024; Ayupov and Chirkova, 2022; Wang et al., 2025b; Yang et al., 2024a; Liu et al., 2024a), (Wang et al., 2025a) CMG(Fan et al., 2024; Liu et al., 2024b), CRF(Hajipour et al., 2024), CRG(Fan et al., 2024), CSum(Liu et al., 2023b; Ayupov and Chirkova, 2022), CTrans(Liu et al., 2023b; Ayupov and Chirkova, 2022) (Wang et al., 2025a), JITCU(Fan et al., 2024), PBT(Wang et al., 2025a) CCD(Liu et al., 2023b; Ayupov and Chirkova, 2022), CR(Yu et al., 2024), DD(Liu et al., 2024b, 2023b), HFP(Wang et al., 2025a)
QLoRA CComp, CTrans(Liu et al., 2024a); CG, UTG (Liu et al., 2024a); CSum(Afrin et al., 2025; Liu et al., 2024a) $\times$
AdaLoRA APR(Huang et al., 2025) $\times$
$F$-LoRA CG, CSum, CTrans(Ayupov and Chirkova, 2022) CCD(Ayupov and Chirkova, 2022)
Adapter-based Base Adapter CG(Chen et al., 2023b; Ayupov and Chirkova, 2022), CMG(Liu et al., 2024b), CRF(Chen et al., 2023b), CSum(Chen et al., 2023b; Liu et al., 2023b; Ayupov and Chirkova, 2022; Wang et al., 2023a), CTrans(Chen et al., 2023b; Liu et al., 2023b; Ayupov and Chirkova, 2022) CCD(Chen et al., 2023b; Liu et al., 2023b; Ayupov and Chirkova, 2022), CS(Wang et al., 2023a), DD(Chen et al., 2023b; Liu et al., 2024b, 2023b)
L-Adapter $\times$ CCD, CT(Saberi et al., 2024), CT(Goel et al., 2022)
T-Adapter CSum(Saberi et al., 2024) CCD(Saberi et al., 2024)
Prompt-based Prompt Tuning CRF(Peng et al., 2024), CSum, CTrans (Wang et al., 2023c), CSum, CTrans (Liu et al., 2023b), CSum, CTrans (Wang et al., 2022b), MNR(Zhu et al., 2023) CCD(Liu et al., 2023b), CS(Wang et al., 2023c), DD(Liu et al., 2023b; Wang et al., 2023c, 2022b), MNCC(Zhu et al., 2023)
Prefix Tuning APR(Li et al., 2024b), CG(Choi and Lee, 2023), CMG, CRG(Fan et al., 2024), CSum(Xie et al., 2023; Liu et al., 2023b; Choi and Lee, 2023), CTrans(Liu et al., 2023b; Choi and Lee, 2023), JITCU(Fan et al., 2024), UTG(Zlotchevski et al., 2022) CCD, DD(Liu et al., 2023b)
P-Tuning APR(Li et al., 2024b), CG, CRF, CSum, CTrans(Chen et al., 2023b) CCD, DD(Chen et al., 2023b)
Pass-Tuning CG, CRF, CSum, CTrans(Chen et al., 2023b) CCD, DD(Chen et al., 2023b)
(IA)3 APR(Li et al., 2024b; Huang et al., 2025) $\times$
Hybrid/Others BitFit CG, CRF, CSum, CTrans(Chen et al., 2023b) CCD, DD(Chen et al., 2023b)
FF-LoRA + AT CG, CSum, CTrans(Ayupov and Chirkova, 2022) CCD(Ayupov and Chirkova, 2022)
L-Adapter + T-Adapter $\times$ CCD(Goel et al., 2022)
L-Adapter + NER-Adapter + AdapterFusion CRF, CSum(Saberi and Fard, 2023) $\times$
Telly-K CComp, CG, CSum(Shi et al., 2023) CCD, CS(Shi et al., 2023)
MAM CSum, CTrans(Liu et al., 2023b) CCD, DD(Liu et al., 2023b)
Fig. 5 illustrates the distribution of PEFT sub-methods across SE tasks. Frequent methods like Base LoRA, Base Adapter, and Prompt Tuning are widely applied to generative tasks such as Code Generation and Summarization. This heatmap complements Table 2 highlighting the task-specific popularity of each PEFT technique.
Figure 5. Heatmap of PEFT Methods Across Software Engineering Tasks. The numbers represent the count of studies that utilized a particular PEFT method for a given task.
4.2.3. Deep Dive: How Top PEFT Methods Are Applied Across Software Engineering Tasks.
The configuration of PEFT methods often varies significantly across software engineering tasks due to differences in task structure, data characteristics, and evaluation goals. As such, in our systematic literature review, we conducted a targeted analysis to uncover how specific PEFT methods are adapted and tuned for different SE applications. This task-sensitive variability in implementation suggests that a one-size-fits-all approach to PEFT may not be effective across the diverse landscape of SE tasks.
While adapter-based methods are among the most frequently used in the literature, we exclude them from this deep dive to maintain focus on methods for which the reviewed studies consistently report configuration details. This choice allows us to conduct a more granular and meaningful comparison across tasks and studies. For each selected PEFT method, we examine how it has been applied in practice by analyzing its usage across multiple SE tasks. Our goal is to surface trends in implementation strategies, training setups, and adaptation mechanisms that researchers have employed to optimize these methods for their specific needs.
This deep dive not only highlights methodological differences but also reveals emerging best practices and potential gaps in reporting standards. In the following tables, we provide a detailed summary of configurations for each method. We prioritize studies that offer transparent accounts of their fine-tuning procedures, including architectural choices, learning rates, layer selection strategies, and task-specific adjustments.
For compactness, the SE Task column uses abbreviated task names, which are fully listed in Table 1.
LoRA: Among all PEFT techniques, LoRA emerges as the most widely adopted approach across software engineering tasks, being applied in 36 distinct SE task instances. When including its hybrid variant, FF-LoRA combined with Adapter Tuning, the total number of applications extends to 40 SE task instances. To uncover how LoRA is typically configured, we present a synthesis of its use across both generative and non-generative SE tasks in Table 3. The table captures commonly used rank values, target layers (e.g., Attention, Feed-Forward), and the proportion of trainable parameters. By contrasting configurations across task categories, we identify trends that reflect LoRA’s adaptability and efficiency in varied contexts.
Table 3. Application of LoRA Across Software Engineering Tasks with Configuration Details.
Category Study SE Task Rank (r) Alpha ($\alpha$) Target Layers Trainable Params (%)
Generative Li et al. (Li et al., 2024b) APR 32 16 Attention 0.13-0.25% (varying by model)
Liu et al. (Liu et al., 2024c) APR 8 $\times$ $\times$ $\times$
Huang et al. (Huang et al., 2025) APR 16 32 Attention + Feed-Forward (InCoder-6B), Attention $\times$
Wang et al. (Wang et al., 2025b) CG 8 16 $\times$ 3.86%
Hajipour et al. (Hajipour et al., 2024) CG, CRF (16,16) 32 $\times$ <<1%
Fan et al. (Fan et al., 2024) CMG, CRF, JITCU 8 16 Attention $\times$
Liu et al. (Liu et al., 2024b) CMG, DD $\times$ $\times$ $\times$ 0.40%
Non-Generative Yu et al. (Yu et al., 2024) CR 8 $\times$ $\times$ 0.05%
Ayupov et al. (Ayupov and Chirkova, 2022) CCD 1, 4, 16 $\times$ Attention 0.7–1.9%
Ayupov et al. (Ayupov and Chirkova, 2022) CG 256, 512 $\times$ Attention Up to 74%
Ayupov et al. (Ayupov and Chirkova, 2022) CSum 2, 4, 8, 16 $\times$ Attention 0.1–2.3%
In generative tasks, LoRA is commonly applied to either the Attention layers alone or both the Attention and Feed-Forward layers, with typical rank values such as 8 or 16. These configurations strike a balance between performance and efficiency, achieving competitive results with minimal additional parameter overhead.
In contrast, non-generative tasks tend to prioritize efficiency more strongly, often utilizing lower rank values (e.g.,, 1, 4, or 8), which keep the number of trainable parameters well below 1% of the base model’s total.
A distinctive pattern emerges in non-generative contexts, where much higher ranks (256 or 512) are commonly used, particularly with FF-LoRA implementations that limit adaptation to Feed-Forward layers only. This phenomenon likely stems from the distinctive requirements of non-generative software engineering tasks like defect detection, classification, and code clone identification, which rely more heavily on discriminative learning and detailed feature extraction. The higher ranks may serve as a compensatory mechanism, enhancing expressiveness and capacity within FF-LoRA’s constrained adaptation scope.
Prompt Tuning: Among the PEFT techniques explored in this review, Prompt Tuning emerges as a versatile strategy, applied in 14 studies across both generative and non-generative software engineering tasks. It includes both hard prompts, which are manually crafted instructions, and soft prompts, which are learnable embeddings prepended to input sequences. Table 4 summarizes how Prompt Tuning has been tailored for different software engineering tasks. In particular, it highlights the types of prompts used (hard, soft, or both), the backbone models they were applied to, and how the prompts were adapted to match task-specific requirements. These variations demonstrate the flexibility of Prompt Tuning in aligning with diverse task goals and model architectures.
Table 4. Application of Prompt Tuning Across Software Engineering Tasks with Configuration Details.
Category Study SE Task Prompt Type Model Used Task-Specific Adaptation
Generative Zhu et al. (Zhu et al., 2023) MNR Hard CodeT5 Prompts guide naming logic and context awareness; output limited to 16 tokens for method names.
Wang et al. (Wang et al., 2023c) CSum Hard + Soft CodeT5-small, CodeT5-base Natural language instruction prepended (e.g.,, “Generate comment…”); soft prompt variants also tested.
Wang et al. (Wang et al., 2023c) CTrans Hard + Soft CodeT5-small, CodeT5-base Prompt specifies the translation task using structured natural language (e.g., “Translate [X] to [LANG] [Z]”).
Wang et al. (Wang et al., 2022b) CSum Hard $\rightarrow$ Soft CodeT5-small, CodeT5-base Replaced natural language instructions with learnable virtual tokens (soft prompts).
Peng et al. (Peng et al., 2024) CRF Domain-Aware Hard CodeT5 Domain-specific templates created from AST edits using <extra_id_X> masked tokens.
Non-Generative Zhu et al. (Zhu et al., 2023) MNCC Hard CodeT5 Prompts embed method usage and naming patterns to classify naming consistency.
Wang et al. (Wang et al., 2023c) $D$ Hard + Soft CodeBERT, CodeT5-small, CodeT5-base Cloze-style templates with verbalizers mapping to “defective” or “clean”.
Wang et al. (Wang et al., 2022b) $D$ Hard $\rightarrow$ Soft CodeBERT, CodeT5-small, CodeT5-base Symbolic prompt labels like + and -.
Wang et al. (Wang et al., 2023c) CS Hard + Soft CodeBERT Language-specific template for retrieval scoring.
Table 4 illustrates how Prompt Tuning has been flexibly applied across a variety of software engineering tasks, showcasing both its adaptability and efficiency. Notably, Wang et al. (Wang et al., 2023c) experimented with Prompt Tuning across four distinct SE tasks—Code Summarization (CSum), Code Translation (CTrans), Code Search (CS), and Defect Detection (DD)—and adopted different prompting strategies tailored to the specific characteristics of each task. For generative tasks like Code Summarization and Code Translation, they used a combination of hard and soft prompts. In Code Summarization, natural language instructions (e.g.,, “Generate comment…”) were prepended to the input to steer the generation process, and later replaced with learnable virtual tokens to test soft prompt effectiveness. For Code Translation, the prompts adopted structured natural language templates (e.g., “Translate [X] to [LANG] [Z]”), clearly specifying the translation goal.
In non-generative tasks, such as Code Search and Defect Detection, prompting strategies were adjusted accordingly. For Code Search, Wang et al. employed language-specific hard templates optimized for retrieval scoring. For Defect Detection, they explored both symbolic labels (e.g., + and -) and cloze-style templates with verbalizers mapping to labels like “defective” or “clean.”
These configurations underscore how prompt design is carefully aligned with task functionality—leveraging natural language guidance for text generation and structured or symbolic cues for classification and retrieval. Moreover, studies such as Peng et al. (Peng et al., 2024) extended this adaptability further by introducing domain-aware hard prompts derived from abstract syntax tree (AST) edits, reinforcing Prompt Tuning’s potential in context-sensitive SE tasks like code repair.
Overall, Prompt Tuning proves to be a highly versatile PEFT method, enabling task-specific customization with minimal parameter updates. Its support for both static (hard) and dynamic (soft) prompt forms allows for fine-grained control over model behavior, making it a compelling alternative to full fine-tuning in diverse software engineering scenarios.
Prefix Tuning. Through prefix vectors across all layers of the network, Prefix Tuning is employed in 13 studies across both generative and non-generative SE tasks. It adapts models by pre-pending trainable vectors, referred to as prefix tokens, to the input or intermediate layers, allowing task-specific customization without updating model weights. This approach enables parameter-efficient adaptation, particularly in resource-constrained scenarios. Table 5 presents key configuration details, including the length of prefix tokens, projection dimensions, and task-specific adaptations such as domain-informed initialization, integration with complementary PEFT methods such as LoRA, and architecture-aware tuning strategies.
Table 5. Application of Prefix Tuning Across Software Engineering Tasks with Configuration Details.
Category Study SE Task Prefix Length Projection Dim (m) Task-Specific Adaptation
Generative Li et al. (Li et al., 2024b) APR 100 tokens (all layers) 256 Standard prefix tuning across layers via Hugging Face PEFT.
Choi et al. (Choi and Lee, 2023) CG 250 tokens $\times$ Input-dependent prompts with multi-word initialization.
Choi et al. (Choi and Lee, 2023) CSum 200 tokens $\times$ Corpus-specific prefix embeddings with task-tuned length.
Choi et al. (Choi and Lee, 2023) CTrans 100 tokens $\times$ Shorter prefix used to guide structural mappings in translation.
Xie et al. (Xie et al., 2023) CSum $\times$ 512 Frozen transformer; trained final linear layer only.
Fan et al. (Fan et al., 2024) CMG, CRG, JITCU 8 vectors $\times$ Prefix tuning was used alongside LoRA (rank=8) to enable dual adaptation of attention layers and prompt control.
Non-Generative Liu et al. (Liu et al., 2023b) CCD, DD 16 tokens 400 Integrated into MAM setup; used rank=16, 8 heads, projection dim=200 for prefix mixing.
Zlotchevski et al. (Zlotchevski et al., 2022) UTG 200 tokens $\times$ Domain-initialized tokens from frequent repository terms.
As shown in Table 5, Prefix Tuning has been applied across diverse SE tasks with considerable variation in configuration. In generative tasks, notable is the work by Choi et al. (Choi and Lee, 2023) in which Prefix Tuning has been explored for different scenarios, including Code Generation, Code Summarization, and Code Translation. For each specific task, the prefix lengths were adjusted to encode the naturally-occurring semantics of the task itself.
Li et al. (Li et al., 2024b) used a 100-token prefix across all layers with a projection dimension of 256 for Automated Program Repair, following the standard Hugging Face PEFT implementation [^4] Refers to the PrefixTuningConfig and PrefixEncoder modules in the Hugging Face peft library (https://github.com/huggingface/peft), which support prefix token insertion with optional projection layers to reduce dimensionality and improve efficiency..
Continuing in this direction, Xie et al. (Xie et al., 2023) employed Prefix Tuning by freezing all transformer layers and training only the final linear layer, using a projection dimension of 512. In the context of non-generative tasks, Liu et al. (Liu et al., 2023b) integrated Prefix Tuning into a Mix-And-Match Adapter (MAM) framework for Code Change Detection and Defect Detection, configuring it with 16-token prefixes and a projection dimension of 400. These variations demonstrate the adaptability of Prefix Tuning across SE tasks, allowing targeted task conditioning with minimal parameter updates.
By tuning only a small set of prefix parameters while keeping the underlying model weights fixed, Prefix Tuning facilitates efficient fine-tuning with lower computational overhead, making it particularly suitable for resource-constrained or multi-task settings.
{iconbox}
Key Findings (RQ2)[Uncaptioned image]
Our analysis reveals consistent trends in the use of model architectures and PEFT methods across software engineering tasks. Encoder-decoder architectures are the most commonly used (29 studies), followed by encoder-only (22) and decoder-only (19). Among base models, CodeT5 appears most frequently (19 studies), particularly in generation tasks, followed by GraphCodeBERT (10), CodeBERT and PLBART (9 each), and UniXcoder (8). Variants such as CodeT5-base (10 studies) and the CodeLlama series (7B, 13B, 34B) show increasing adoption, along with emerging large-scale models like StarCoder-15B, DeepSeek-Coder-33B, and LLaMA-2-70B. Base LoRA is the most widely used PEFT technique (29 studies), especially in code generation and program repair. Prompt-based strategies, such as Prompt Tuning and Prefix Tuning, are widely adopted—especially in tasks like Code Summarization and Code Translation. Our in-depth analysis of three PEFT methods highlights that LoRA is often preferred for its strong balance between performance and parameter efficiency, while Prompt and Prefix Tuning are valued for their flexible input-conditioning capabilities. Notably, some studies explore hybrid configurations, combining Prefix Tuning with LoRA, suggesting early experimentation with multi-strategy adaptation. Despite these advances, configuration details are not consistently reported across the literature, particularly for adapter-based approaches, limiting reproducibility and comparative analysis. Architecture-task alignment analysis reveals that decoder-only models appeared exclusively in generative tasks, suggesting fundamental architectural requirements for different SE task categories and indicating a strong correlation between model design and task functionality.
{iconbox}
Future Directions (RQ2)[Uncaptioned image]
Future research should continue leveraging larger models such as CodeLlama-13B/34B, StarCoder-15B, and DeepSeek-Coder-33B, which have shown growing adoption across tasks but still offer untapped potential for broader PEFT applications. There is also a need to examine architecture-task alignment more systematically, as decoder-only models show strong performance in generative tasks like Code Generation and Program Repair. Although LoRA is the most widely used PEFT method, future work should explore combining it with other strategies like Adapters or Prefix Tuning to improve adaptability. The success of quantization-aware techniques such as QLoRA also suggests opportunities for scaling large models efficiently. Finally, developing task-agnostic PEFT frameworks that generalize across software engineering tasks could enhance both reusability and performance in real-world scenarios. Future work may also explore the design space of hybrid PEFT combinations (e.g., LoRA + Prefix) and standardized benchmarking protocols to evaluate PEFT efficacy across diverse software engineering tasks.
4.3. RQ3: How do PEFT methods impact task performance and training efficiency in software engineering tasks when compared against full fine-tuning?
Having outlined the distribution of PEFT methods across various SE tasks and model architectures in Section 4.2, we now examine how these methods perform in practice. In this section, we examine empirical findings from studies that benchmark PEFT techniques against valid baselines—specifically, the same base model fine-tuned via full-parameter updates-across both generative and non-generative software engineering tasks. For each study, we select and report only the best-performing PEFT method to highlight the most effective adaptation strategy per paper. Our evaluation focuses on two key dimensions: (i) task-specific performance–assessed by using BLEU, METEOR, Pass@k, Accuracy, F1 etc.; (ii) training-time and/or memory efficiency, when explicitly reported.
To support this analysis, Table 6 presents an overview of PEFT techniques and their corresponding sub-variants, summarizing their relative effectiveness and efficiency across different categories of software engineering tasks. The table uses symbolic indicators to denote whether each PEFT method demonstrated performance improvements, degradations, or parity relative to its full fine-tuning counterpart. This analysis of performance and resource usage provides insight into which PEFT strategies offer competitive results at reduced computational cost, and under which task-specific conditions, they are most effective.
Table 6. Summary of Best-Performing PEFT Methods Across Software Engineering Tasks. We report both performance impact ( P= improved, P= declined, P= similar) and training/memory efficiency ( E= more efficient, E= less efficient, E= similar), based on comparisons with corresponding full fine-tuning baselines. If only performance is reported (e.g., P), it indicates that efficiency was not addressed in the original study. Each entry represents the best-performing PEFT variant identified in the respective study for the given task and base model. Tasks are grouped into generative and non-generative categories for clarity.
Main Method Sub-Method Generative Tasks Non-Generative Tasks
LoRA-based Base LoRA APR(Huang et al., 2025) P ECG(Wang et al., 2025b) P E, CG(Fakih et al., 2024) P E, CG(Hajipour et al., 2024) P, CG(Yang et al., 2024a) P, CG(Wang et al., 2025a) P, CComp(Wang et al., 2025a) P, CRF(Hajipour et al., 2024) P, CSum(Ayupov and Chirkova, 2022) P, CTrans(Wang et al., 2025a) P, PBT(Wang et al., 2025a) P CCD(Ayupov and Chirkova, 2022) P, CR(Yu et al., 2024) P, DD(Liu et al., 2024b) P, HFP(Wang et al., 2025a) P
QLoRA CComp(Liu et al., 2024a) P E, CG(Liu et al., 2024a) P E, CSum(Afrin et al., 2025) P E $\times$
AdaLoRA $\times$ $\times$
$F$-LoRA $\times$ $\times$
Adapter-based Base Adapter CMG(Liu et al., 2024b) P E, CSum(Wang et al., 2023a) P, CTrans(Chen et al., 2023b) P CCD(Chen et al., 2023b) P, CCD(Liu et al., 2023b) P, CS(Wang et al., 2023a) P
L-Adapter $\times$ CT(Saberi et al., 2024) P E
T-Adapter CSum(Saberi et al., 2024) P E $\times$
Prompt-based Prompt Tuning CSum(Wang et al., 2023c) P, CSum(Wang et al., 2022b) P, CTrans(Wang et al., 2023c) P, CTrans(Wang et al., 2022b) P CS(Wang et al., 2023c) P, DD(Wang et al., 2023c) P, DD(Wang et al., 2022b) P
Prefix Tuning $\times$ $\times$
P-Tuning $\times$ $\times$
Pass-Tuning CG(Chen et al., 2023b) P, CRF(Chen et al., 2023b) P, CSum(Chen et al., 2023b) P $\times$
(IA)3 APR(Li et al., 2024b) P E $\times$
Hybrid/Others BitFit $\times$ $\times$
FF-LoRA + AT CTrans(Ayupov and Chirkova, 2022) P CCD(Ayupov and Chirkova, 2022) P
L-Adapter + T-Adapter $\times$ CCD(Saberi et al., 2024) P E
L-Adapter + NER-Adapter + AdapterFusion CSum(Saberi and Fard, 2023) P E, CRF(Saberi and Fard, 2023) P E $\times$
Telly-K CComp(Shi et al., 2023) P E, CG(Shi et al., 2023) P E, CSum(Shi et al., 2023) P E CCD(Shi et al., 2023) P E, CS(Shi et al., 2023) P E
MAM CSum(Liu et al., 2023b) P, CTrans(Liu et al., 2023b) P DD(Liu et al., 2023b) P
As anticipated, our analysis focuses exclusively on studies that benchmark PEFT methods against full fine-tuning, enabling controlled evaluations grounded in clearly reported performance and efficiency metrics. This selection criterion led to the inclusion of 20 studies in our evaluation. At the same time, it highlights a notable gap in the literature: $\sim$29% of the reviewed works do not include any direct comparison with full fine-tuning model, making it difficult to assess the practical trade-offs of PEFT methods in those cases.
In generative tasks, the application of PEFT methods demonstrates to produce results on par to full model’s parameters fine-tuning or surprisingly even better in both accuracy and efficiency. Based on the evidence from the 34 generative task entries in Table 6, ~72% report performance improvements ( P), and over half (~54%) also report gains in training or memory efficiency ( E). These trends are most pronounced in Code Summarization, Code Generation, and Code Refinement, which collectively represent the most frequently studied generative SE tasks in PEFT evaluations.
Among the various strategies, LoRA-based methods—particularly Base LoRA and QLoRA—stand out as the most widely adopted and effective. Base LoRA is applied across four distinct generative tasks: i.e., Automated Program Repair, Code Generation, Code Refinement, and Code Summarization. All cases report improved task performance ( P), with efficiency gains ( E) observed in APR and one Code Generation study, while one Code Generation study reports a slight drop in efficiency ( E). Similarly, QLoRA shows high effectiveness in Code Generation, Code Completion, and Code Summarization, consistently yielding performance and efficiency gains ( P E), further solidifying the role of LoRA-based methods as leading choices for token-level generation tasks. Recent comprehensive evaluations by Wang et al. (Wang et al., 2025a) provide additional empirical validation of these performance patterns while revealing important limitations in domain-specific scenarios. Their systematic comparison across five code intelligence tasks demonstrates that while PEFT methods like LoRA can achieve competitive performance on public datasets, substantial performance gaps emerge when applied to private data contexts. Their findings reveal mixed effectiveness of PEFT methods, with competitive performance on general-domain tasks but notable performance degradation in scenarios involving private codebases and specialized domain knowledge, where the distribution shift from pre-training data was substantial. This evidence reinforces our systematic findings that while LoRA demonstrates overall effectiveness across diverse applications, significant challenges remain in complex scenarios where substantial domain adaptation and knowledge acquisition beyond the model’s pre-training data are required. Adapter-based methods show more varied results. In particular, Base Adapter achieves performance improvements in Code Summarization and Commit Message Generation—with the latter performing on par to the full fine-tuned model when evaluated for generating code summaries ( P), but gaining in efficiency ( E). However, its application to Code Translation results in a performance decline ( P), a trend mirrored by the hybrid FF-LoRA + AT and MAM approaches. These declines across three different studies indicate that Code Translation may be less amenable to parameter-efficient strategies, likely due to its higher sensitivity to sequence-level alignment and cross-language mapping.
On the other hand, Prompt-based methods yield generally strong results in generative contexts. Prompt Tuning is successfully applied to Code Summarization and Code Translation, achieving consistent performance gains ( P). Likewise, Pass-Tuning and (IA)3 also yield improvements in specific tasks—Code Refinement, Code Summarization, and Automated Program Repair, respectively—with (IA)3 additionally reporting efficiency gains ( E). However, Pass-Tuning shows a performance decline ( P) in one Code Generation instance, suggesting that tuning strategies may not be uniformly effective across all task types or data distributions.
Hybrid and other PEFT methods such as Telly-K, L-Adapter + NER-Adapter + AdapterFusion demonstrate promising performance, particularly in Code Summarization and Code Refinement. Notably, Telly-K reports performance and efficiency gains ( P E) in three generative tasks—Code Generation, Code Summarization, and Code Completion—suggesting that selective fine-tuning of higher transformer layers can effectively retain model expressiveness while reducing overhead. In contrast, MAM shows performance degradation ( P) in both Code Summarization and Code Translation. This may be attributed to limitations in modular adaptation when the task requires transforming data from one domain and projecting it onto a different target domain (e.g., translate code into natural language).
In non-generative tasks, PEFT methods demonstrate a more nuanced performance profile. Based on the 16 non-generative task entries in Table 6, approximately $\sim$62% report improved task performance ( P), while around $\sim$50% show gains in efficiency ( E). This reflects moderate but meaningful benefits of PEFT in retrieval and classification SE tasks. The most frequent task is Code Clone Detection, which shows mixed results. In this case, we count three entries where improved or neutral performance, three show performance degradation ( P), indicating sensitivity to semantic precision and token similarity under compressed training regimes. In contrast, Defect Detection tasks demonstrate consistent improvements in performance, particularly when Prompt Tuning or LoRA is used.
Among PEFT strategies, LoRA-based methods (e.g., Base LoRA) report success in Defect Detection and Code Review, with consistent performance improvements ( P). However, its application to Code Clone Detection shows performance degradation ( P), suggesting potential limitations in handling fine-grained syntactic variance. Adapter-based methods yield mixed outcomes. For instance, Base Adapter leads to performance improvements in Code Search, but its results in Code Clone Detection are not so evident—improving in one study and degrading in another. The application of L-Adapter to the Cloze Test task yields improvements in both performance and efficiency ( P E), indicating a promising outcome for structured classification tasks. Prompt-based methods, especially Prompt Tuning, yield consistent performance gains in Defect Detection and Code Search, aligning well with their token-efficient query-focused design. Notably, these methods are not applied to Code Clone Detection or Code Review in the current set of studies, leaving room for future exploration. Finally, hybrid and selective tuning methods, such as Telly-K and L-Adapter + T-Adapter, demonstrate performance gains in tasks like Code Clone Detection and Code Search ( P E).
On a different, yet related note, our analysis of Table 6 reveals a notable trend in the literature: researchers are increasingly employing PEFT methods not only to improve fine-tuning efficiency but also to enhance model performance on the target task. While this may initially appear counterintuitive—given that PEFT techniques are typically associated with reducing computational overhead—our systematic review identified at least 7 studies across generative tasks where the primary motivation was performance optimization. In these studies, PEFT techniques such as LoRA, QLoRA, Prompt Tuning, and Adapter Fusion were adopted with the explicit goal of maximizing task-specific accuracy or generalization, rather than solely minimizing resource usage. This dual role of PEFT—as a strategy for achieving both computational efficiency and performance improvements, or the latter alone–communicates its versatility and growing relevance in generative software engineering tasks. On an interesting note, the recent paper by Wang et al. (Wang et al., 2025a) provides valuable insights into the effectiveness of PEFT methods for software engineering across both private (non–open-source) and public (open-source) application domains. Their findings show that, while PEFT methods perform competitively on general-domain tasks, performance declines markedly in scenarios requiring substantial adaptation to previously unseen knowledge. This suggests that PEFT’s effectiveness may be inherently constrained by the degree of domain-specific adaptation needed. However, this outcome may also be influenced by the predominant reliance of large code models on open-source training data–typically sourced from public GitHub repositories. As a result, the observed performance drop could partly stem from a data distribution shift caused by applying PEFT across two distinct distributions: the pre-training dataset and the domain-specific data used for PEFT-optimized fine-tuning. As the demand for deploying large language models in real-world development workflows continues to rise, PEFT emerges as a practical and scalable alternative to full fine-tuning, offering significant gains in both accuracy and adaptability across diverse code generation scenarios.
{iconbox}
Key Findings (RQ3)[Uncaptioned image]
Our analysis reveals significant disparities in PEFT method coverage across software engineering tasks. While LoRA-based and adapter-based demonstrate consistent effectiveness (both in generative and non-generative tasks), we note a striking absence of empirical evaluations for several established PEFT methods—including AdaLoRA, Prefix Tuning, P-Tuning, and BitFit. The current evidence strongly favors LoRA variants for Code Generation, whereas non-generative tasks show more varied outcomes, with adapter-based emerging as the most used approach. Performance analysis reveals that $\sim$72.7% of generative tasks report performance improvements compared to $\sim$62.5% of non-generative tasks, demonstrating fundamental differences in PEFT applicability across task categories. Efficiency-performance correlation analysis shows dual benefits, with over half of generative tasks achieving efficiency gains alongside performance improvements, while around $\sim$25% of non-generative tasks demonstrate efficiency gains. Critical analysis of Base LoRA shows $\sim$60% success rates against full fine-tuning baselines, challenging conventional assumptions about parameter efficiency trade-offs and suggesting that targeted parameter updates may be more effective than comprehensive fine-tuning for specific software engineering contexts.
{iconbox}
Future Directions (RQ3)[Uncaptioned image]
While current evidence demonstrates the effectiveness of certain PEFT methods (particularly LoRA variants and Adapter-based) for software engineering tasks, the limited evaluation of other promising approaches like AdaLoRA, Prefix-Tuning, P-Tuning, and BitFit leaves their potential largely unexplored. Future work should expand the methodological diversity of PEFT evaluations through standardized benchmarking across different task categories, which would enable more comprehensive comparisons and help identify optimal adaptation strategies for specific SE scenarios. Such efforts could particularly benefit from investigating how understudied methods might address current limitations in parameter efficiency and task performance, while establishing clearer guidelines for method selection based on task characteristics and resource constraints. Additionally, future research must address domain-specific PEFT optimization that accounts for software engineering’s unique characteristics. This enhancement should include recommendations for adapting PEFT methods to preserve code structure and formatting requirements, providing evidence-based guidance that distinguishes between methods suitable for structured code generation versus natural language synthesis in software engineering contexts. Research should emphasize the importance of code-specific evaluation metrics such as syntactic correctness and compilability alongside traditional performance measures, while investigating PEFT methods specifically designed for code’s hierarchical and syntactic properties. This comprehensive approach will provide the domain-aware guidance necessary for effective PEFT adaptation in software engineering, recognizing that methods optimized for general natural language processing may require significant modification to handle the structured, syntax-dependent nature of code and software engineering artifacts.
5. Practical Implications and Strategic Recommendations
Building on our systematic analysis of PEFT applications in software engineering, we synthesize our findings into actionable insights across several critical dimensions: methodological strategy, resource optimization.
Methodological Strategy Implications: The documented concentration in synthesis-based applications suggests that organizations should prioritize PEFT investment in code generation and summarization infrastructure, as these represent the most mature and validated use cases. The exclusive association of decoder-only architectures with generative tasks provides clear architectural selection criteria for practitioners. Organizations implementing code generation systems should prioritize decoder-only models, while those requiring mixed analytical and generative capabilities should consider encoder-decoder architectures. The demonstrated reliability of Base LoRA across diverse applications establishes it as a primary choice for initial PEFT implementations, particularly when performance parity with full fine-tuning is critical.
Resource Optimization Implications: The superior performance-efficiency correlation in generative tasks suggests that organizations should prioritize PEFT adoption for generative applications to maximize return on computational investment. The lower but still substantial benefits in non-generative tasks indicate that PEFT adoption should follow a staged approach, beginning with generative applications before expanding to analytical tasks. The evidence for performance enhancement beyond traditional fine-tuning suggests that PEFT should not be viewed merely as a resource-saving compromise, but as a potentially superior adaptation strategy. This paradigm shift has significant implications for model development budgets and computational resource allocation in software engineering organizations. The current state of PEFT research demonstrates remarkable community progress, with widespread availability of optimized implementations through platforms like HuggingFace PEFT, facilitating practical adoption and reproducibility across software engineering applications.
Our findings demonstrate that PEFT optimization methods, including both parameter-efficient approaches like LoRA and prompt-based methods like Prompt Tuning and Prefix Tuning, often match or exceed full fine-tuning performance while requiring significantly fewer computational resources. This enables researchers to establish strong performance baselines through PEFT evaluation without requiring computationally expensive full model fine-tuning, promoting both research efficiency and environmental sustainability in large language model adaptation for software engineering tasks.
6. Future Directions for PEFT Applications in Software Engineering
Building on the findings of our systematic literature review, we identify several future research directions to advance the application of PEFT techniques in SE-related activities. These directions synthesize insights across generative and non-generative tasks, model architectures, PEFT strategies, and observed evaluation gaps. We organize them into six actionable steps designed to guide both researchers and practitioners in the field.
Step 1: Identify Underexplored and Emerging SE Tasks: While PEFT has seen wide application in tasks like Code Summarization and Generation, Several SE tasks remain underrepresented in the current literature. These include, but are not limited to, Software Testing, Documentation Generation, Code Quality Assessment, API Recommendation, Automated Refactoring, and Requirements Analysis. As PEFT research matures, there is ample opportunity to expand into additional task categories and workflows that are critical to real–world software development. Future work should target these areas, especially those involving complex or cross–functional requirements. Cross–task adaptation also emerges as a promising direction, where a single PEFT configuration generalizes across multiple tasks, reflecting real–world development scenarios more effectively.
What to report: Clearly define the SE task and justify its importance. If proposing a novel or emerging task, document its scope, challenges, and dataset preparation methodology to aid reproducibility. Discuss why PEFT is suitable and note potential for cross–task generalization.
Step 2: Align Base Models with Task Requirements: Our findings emphasize the prevalence of encoder–decoder architectures, with CodeT5 and CodeLlama variants showing dominant usage, particularly in generative tasks. Larger models like StarCoder–15B and DeepSeek–Coder–33B are gaining adoption but remain underutilized. There is also an opportunity to experiment with hybrid or task–agnostic model architectures that can support a range of SE tasks under a unified PEFT configuration.
What to report: Justify model selection based on task characteristics. Include details on architecture type (encoder–only, decoder–only, encoder–decoder), model scale, and pre–training corpus. If hybrid or ensemble models are used, explain their complementary advantages.
Step 3: Expand and Diversify PEFT Techniques: LoRA has emerged as the most adopted PEFT method (used in 36 studies), especially for generative tasks. Adapter-based methods dominate in non-generative settings. However, many promising PEFT approaches–such as AdaLoRA, BitFit, Prefix Tuning, and P-Tuning remain significantly under-evaluated. Additionally, advanced LoRA variants including LoRA-pro and LoRA-GA remain unexplored in software engineering contexts, despite their potential for enhanced parameter efficiency and gradient-aware optimization, respectively. Given LoRA’s demonstrated $\sim$77% success rate against full fine-tuning baselines in our analysis, these sophisticated variants could potentially address performance limitations observed in certain software engineering tasks, particularly where traditional PEFT methods showed mixed results. Future research should not only expand coverage to these methods but also explore hybrid strategies (e.g., LoRA + Prefix Tuning) and configurations tailored to task type. Furthermore, quantization-aware variants such as QLoRA demonstrate the feasibility of scaling large models efficiently and warrant deeper investigation.
What to report: Detail the PEFT methods used, including training configurations, memory reduction techniques (e.g., 8–bit/4–bit quantization), and note whether any hybrid or novel PEFT combinations were tested. Clearly state why a particular method was chosen for a given task and architecture.
Step 4: Evaluate Adaptability, Efficiency, and Generalizability: PEFT research in SE should go beyond accuracy metrics and include broader evaluations—covering generalization across tasks and programming languages, memory utilization during training/inference, and performance in low-resource settings. Our review highlights how LoRA, Adapters, and Prompt Tuning excel in cross-domain and low-data scenarios, yet few studies offer comprehensive ablation or scalability analysis. Underexplored methods should also be evaluated in terms of parameter efficiency and hardware suitability (e.g., edge deployment).
What to report: Include evaluation metrics that match the task (e.g., BLEU, pass@1 for generative tasks; F1, MRR, AUC for non–generative tasks). Report memory and latency improvements over full fine–tuning, and include results of generalizability studies, ablations, and cross–task as well as cross–lingual evaluations.
Step 5: Promote Reproducibility and Benchmarking Platforms: Despite progress in tasks like Code Generation and Clone Detection, inconsistency in dataset usage and reporting limits comparability. Adapter–based studies, in particular, often lack detailed configuration descriptions. Future work should adopt standardized datasets and benchmarking protocols, share full implementation details, and follow community practices for reproducibility. Benchmarking diverse PEFT techniques side by side on shared tasks would help identify the most effective and efficient approaches.
What to report: Provide all scripts, datasets, and model checkpoints through public repositories. Document hardware specs, training steps, and evaluation protocols. Follow consistent reporting formats, especially when introducing or comparing multiple PEFT techniques.
Step 6: Address Domain-Specific Knowledge Acquisition Limitations: Emerging evidence reveals a critical gap in current PEFT research–the fundamental challenge of adapting to domain-specific knowledge not present in pre-training data (Wang et al., 2025a). PEFT methods face substantial performance degradation when applied to private codebases or specialized domains, highlighting the limitation of updating only parameter subsets for comprehensive knowledge acquisition. Future research should prioritize developing PEFT variants specifically designed for domain adaptation scenarios, including techniques that can effectively encode new knowledge while maintaining parameter efficiency. This includes exploring progressive adaptation strategies, domain-aware parameter selection methods, and hybrid approaches that combine PEFT efficiency with targeted full parameter updates for critical knowledge domains.
What to report: Clearly distinguish between evaluations on public and private datasets, and, where applicable, report domain shift metrics–such as Jensen–Shannon divergence (Lin, 1991) or Kullback–Leibler divergence (Kullback and Leibler, 1951)–to quantify distribution differences. Complement performance results with an analysis of knowledge acquisition effectiveness, and document the specific domain characteristics and adaptation challenges considered.
Together, these six steps outline a comprehensive roadmap for advancing the application of PEFT in software engineering. By broadening the scope of SE tasks, aligning base models with task demands, diversifying PEFT techniques, expanding evaluation criteria, and promoting reproducibility, future research can address current limitations while laying the groundwork for more robust, efficient, and generalizable fine-tuning solutions. We believe that systematically adopting these practices will not only accelerate progress in the field, but also ensure that PEFT methods are better positioned to meet the diverse and evolving needs of real-world software development.
7. Threats to Validity
While our systematic literature review follows the established guidelines of Kitchenham et al. (Kitchenham et al., 2009), we still have to acknowledge the presence of threats to validity that can hinder our investigation. We discuss these threats in terms of construct, internal, and external validity.
Construct Validity relates to how accurately we identified and categorized core elements such as software engineering tasks, model architectures, and PEFT methods. To mitigate the risk of misclassifications, our approach is grounded in well-established software engineering practices relevant to the field (e.g., Watson et al. (Watson et al., 2022)) and is developed according to the guidelines by Kitchenham et al. (Kitchenham et al., 2009). Additionally, multiple authors conducted independent reviews to ensure the accuracy and completeness of the task and method selection.
We also recognize that some practical uses of PEFT–such as those embedded in open-source repositories–may not be explicitly documented in peer-reviewed publications. While these cases may reflect real-world adoption, they often lack the methodological transparency and empirical evaluation required for systematic analysis. For instance, enabling PEFT through configuration toggles in frameworks like HuggingFace Transformers is rarely accompanied by sufficient detail on how the method was integrated, assessed, or compared against baselines. Including such undocumented or implicit usage would conflict with our review’s goal of synthesizing empirically validated and reproducible studies.
Internal Validity addresses whether our findings accurately represent trends without being biased by our selection or extraction process. Each included study was reviewed by at least two authors, with classification results cross-checked and refined. While this process helped ensure accuracy, potential inconsistencies may still exist due to limited methodological details in some primary studies, particularly regarding adapter-based or hybrid PEFT configurations.
External Validity concerns the generalizability of our results. Although we used broad search strings and included top-tier venues identified through CSRankings, some relevant studies may have been missed due to terminology variation, indexing limitations, or query constraints in platforms like Google Scholar. To address this, we supplemented our search with snowballing and manual inclusion of influential studies. Still, there is a possibility of selection bias and under-representation of work published in non-ranked venues.
8. Data Availability
To support reproducibility and allow for independent verification, we have made our search strings, extracted datasets, and study classifications publicly available in our replication package (rep, 2025).
9. Conclusions
In this article, we conducted a systematic literature review on the application of PEFT techniques in software engineering tasks, guided by Kitchenham’s established methodology (Kitchenham et al., 2009). Through a rigorous multi-phase study selection process, we identified 28 primary studies covering 19 diverse SE tasks, which we grouped into generative and non-generative categories.
The collective findings from the analyzed studies paint an encouraging picture: PEFT techniques have been applied across a wide array of SE tasks and consistently demonstrate the ability to achieve model performance on par with full fine-tuning while dramatically reducing the amount of learned parameters. Across both generative and non-generative tasks, researchers have leveraged PEFT to adapt pre-trained LCMs without incurring significant accuracy loss. In many cases, the best PEFT approaches even marginally outperform traditional fine-tuning baselines on certain metrics, underscoring that careful, efficient adaptation can unlock further performance gains. Moreover, the adoption of PEFT has proven versatile as we found evidence of its success in diverse activities ranging from Code Summarization and Translation to Code Search, Automated Program Repair, and beyond. This breadth of applicability suggests that PEFT methods are not limited to a narrow set of problems, but rather offer a general paradigm for customizing large code models to varied software engineering needs in a resource-conscious manner.
Beyond mapping current trends, this review highlights notable gaps in evaluation rigor and benchmarking consistency, offering a structured set of future directions to advance research in this space.
Overall, the insights from our review solidify the understanding that PEFT is not just a workaround for scarce resources, but a robust optimization paradigm that will likely play a central role in the future of software engineering-focused LCMs, allowing a greener, faster and smarter fine-tuning for the tasks that matter most to practitioners.